### v1

# Non-Expansive Mappings in Two-Time-Scale Stochastic Approximation: Finite-Time Analysis

Abstract
Two-time-scale stochastic approximation is an iterative algorithm used in applications such as optimization, reinforcement learning, and control. Finite-time analysis of these algorithms has primarily focused on fixed point iterations where both time-scales have contractive mappings. In this paper, we study two-time-scale iterations, where the slower time-scale has a non-expansive mapping. For such algorithms, the slower time-scale can be considered a stochastic inexact Krasnoselskii-Mann iteration. We show that the mean square error decays at a rate $\mathcal{O}(1/k^{1/4-\epsilon})$, where $\epsilon>0$ is arbitrarily small. We also show almost sure convergence of iterates to the set of fixed points. We show the applicability of our framework by applying our results to minimax optimization, linear stochastic approximation, and Lagrangian optimization.
keywords:
Two-time-scale stochastic approximation, non-expansive mapping, Krasnoselskii-Mann iteration, finite-time bound, minimax optimization
{MSCcodes}
47H09, 62L20, 47H10, 68W40, 93E35, 90C47
1 Introduction
Stochastic Approximation (SA) is a class of iterative algorithms to find the fixed point of an operator given its noisy realizations [24]. These algorithms have found applications in various areas, including optimization, reinforcement learning [26], communications and stochastic control [2]. Two-time-scale stochastic approximation is a variant of SA, where the iteration is as follows:
$x_{k+1}=x_{k}+\alpha_{k}(f(x_{k},y_{k})-x_{k}+M_{k+1})$
$y_{k+1}=y_{k}+\beta_{k}(g(x_{k},y_{k})-y_{k}+M^{\prime}_{k+1}).$
Here $x_{k}\in\mathbb{R}^{d_{1}}$ and $y_{k}\in\mathbb{R}^{d_{2}}$ are the coupled iterates updated on two separate time-scales, decided by the step sizes $\alpha_{k}$ and $\beta_{k}$. $M_{k+1}$ and $M^{\prime}_{k+1}$ are noise sequences. The iterate $x_{k}$ updates on the faster time-scale, i.e., the step size $\alpha_{k}$ decays slowly. While analyzing the behavior of iterate $x_{k}$, the iterates $y_{k}$ are considered quasi-static or slowly moving. The iterate $y_{k}$ updates on the slower time-scale, i.e., the step size $\beta_{k}$ decays faster. The behavior of iterate $y_{k}$ is often studied in the limit where $x_{k}$ closely tracks $x^{*}(y_{k})$, where $x^{*}(y)$ is the fixed point for $f(\cdot,y)$. The step sizes have to be appropriately chosen for the algorithm to converge.
The majority of analysis for two-time-scale algorithms has occurred under the assumption that the function $f(x,y)$ is contractive in $x$, and the function $g(x^{*}(y),y)$ is contractive [10, 18, 2]. Under these assumptions, the algorithm converges to $x^{*},y^{*}$ where they are the unique solution to the equations: $f(x,y)=x$ and $g(x,y)=y$ [2]. In recent years, there have been works that obtain finite-time bounds for iterates, but they also mainly focus on contractive mappings in both time-scales [10, 18, 16].
In this paper, we consider a non-expansive mapping in the slower iteration, i.e., the mapping $g(x^{*}(y),y)$ is non-expansive. Such two-time-scale iterations arise in applications including minimax optimization [20], Lagrangian optimization [23], and generalized Nash equilibrium problems [5]. The slower time-scale in these iterations can be considered a stochastic inexact Krasnoselskii-Mann (KM) algorithm (an iterative algorithm to find the fixed points of a non-expansive operator). Analysis of these iterations requires novel techniques to deal with the errors due to two-time-scale updates in a stochastic setting with non-expansive maps.
Our main contributions are as follows.
  1. 1.
    We present a finite-time mean square bound for our algorithm, which has contractive and non-expansive mappings in the faster and slower time-scales, respectively. We show that the mean square error decays at a rate $\mathcal{O}(k^{-1/4+\epsilon})$, where $\epsilon>0$ can be arbitrarily small depending on the step size.
  2. 2.
    We also show that the iterates converge to the set of fixed points.
  3. 3.
    We provide a similar result for a projected variant, where the iterates $x_{k}$, updated in the faster time-scale, are projected to a convex and bounded set after each iteration. We show that there exist scenarios where the projection in the faster time-scale leads to the non-expansive mapping (Subsection 5.3).
  4. 4.
    We adapt strongly concave-convex minimax optimization, linear SA, and Lagrangian optimization into our framework.
  5. 5.
    Using similar techniques, we present an additional result where the slower time-scale iteration contains a gradient descent operator for a smooth function. The gradient descent operator is not necessarily non-expansive in this case. In this setting, we show that the minimum, from time 00 to $k$, of the expected norm of the gradient is bounded by $\mathcal{O}(\log(k)/k^{2/5})$.
1.1 Related Work
Applications of SA to reinforcement learning (RL) and optimization have caused a growing interest in obtaining results which bound their finite-time performance. These include mean square error or moment bounds (e.g., [7, 25]) and high probability or concentration bounds (e.g., [6, 8]).
Two-time-scale algorithms with contractive mappings (in both time-scales) play an important role in RL [27], and hence finite time bounds for such algorithms have also been studied in detail. In this setting, the bound is obtained on $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2}]$. A major focus in this field is on linear two-time-scale SA, i.e., where the maps $f(\cdot)$ and $g(\cdot)$ are linear. In [18], an asymptotic convergence rate of $\mathcal{O}(1/k)$ is obtained, while in [16, 17] the result is extended to a finite-time bound. We also study linear two-time-scale SA as a special case. But we study the case where the matrix relevant for the slower time-scale is negative semidefinite instead of Hurwitz or negative definite as assumed in the above works.
Bounds for contractive non-linear two-time-scale SA have also been studied in recent years. In [10], a rate of $\mathcal{O}(1/k^{2/3})$ is obtained for this setting. In [11], a rate of $\mathcal{O}(1/k)$ is obtained but a time-weighted average of iterates is considered.
Two-time-scale SA has also been studied in settings where the mappings are not contractive. In [5], the slower time-scale has a non-expansive mapping, but the result is specific to control of strongly monotone games. In [29, 28], the slower time-scale contains a gradient descent operator $\nabla J(y_{k})$ for a smooth function $J(\cdot)$. The operator $\nabla J(\cdot)$ is not co-coercive in general and is different from our framework. Under this setting, bounds on $\min_{0\leq i\leq k}\mathbb{E}[\|\nabla J(y_{i})\|^{2}]$ are obtained. We give an additional result (Theorem 4.1), where we match their bound of $\mathcal{O}(\log(k)/k^{2/5})$. We work under assumptions different from [29, 28] where they consider Markovian noise but require boundedness of some operators. We also consider the setting where $J(\cdot)$ is convex, which fits into the framework of Theorem 2.1, and obtain a bound on $\mathbb{E}[\|\nabla J(y_{k})\|^{2}]$.
Our slower time-scale can be written as a stochastic inexact KM iteration. Finite-time bounds for inexact KM iterations have been obtained under a variety of settings [19, 9, 21, 12]. In recent years, there have also been works which give bounds for stochastic KM iterations in Banach spaces [3, 4]. Our analysis requires an appropriate combination of errors due to two-time-scale updates with the techniques used for stochastic inexact KM iterations.
1.2 Outline and Notation
The paper is structured as follows: Section 2 sets up and presents the main result. The projected variant is presented in Subsection 2.1. Section 3 contains a proof sketch for our main result. Section 4 presents the framework where the slower time-scale has a gradient descent operator. Section 5 presents the above-listed applications and how they can be incorporated into our framework. Finally, Section 6 concludes the paper and presents some future directions. Appendices A-D contain proofs for our results.
[…]
2 Non-Expansive Mappings in Slower Time-Scale
The primary focus of this paper is on two-time-scale iterations where the faster and slower time-scales are fixed point iterations with contractive and non-expansive mappings, respectively. We present the main results of this paper in this section after setting up the necessary notation and assumptions. Consider the following coupled iterations.
(1) $x_{k+1}=x_{k}+\alpha_{k}(f(x_{k},y_{k})-x_{k}+M_{k+1})$
$y_{k+1}=y_{k}+\beta_{k}(g(x_{k},y_{k})-y_{k}+M^{\prime}_{k+1}).$
Here $x_{k}\in\mathbb{R}^{d_{1}}$ is the iterate updating on the faster time-scale, and $y_{k}\in\mathbb{R}^{d_{2}}$ is the iterate updating on the slower time-scale. $\alpha_{k}$ and $\beta_{k}$ are the respective step sizes that dictate the update speed and are formally defined in Assumption 2. $M_{k+1}\in\mathbb{R}^{d_{1}}$ and $M^{\prime}_{k+1}\in\mathbb{R}^{d_{2}}$ are martingale difference noise sequences (Assumption 2).
We first make the assumption that function $f(x,y)$ is contractive in $x$. {assumption} The function $f:\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{1}}$ is contractive in $x$ for any given $y\in\mathbb{R}^{d_{2}}$, i.e.,
$\|f(x_{1},y)-f(x_{2},y)\|\leq\mu\|x_{1}-x_{2}\|,$
for all $x_{1},x_{2}\in\mathbb{R}^{d_{1}}$ and $y\in\mathbb{R}^{d_{2}}$. Here $\mu<1$ is the contraction factor. By the Banach contraction mapping theorem, the above assumption implies that $f(\cdot,y)$ has a unique fixed point for each $y$. That is, for each $y\in\mathbb{R}^{d_{2}}$, there exists unique $x^{*}(y)$ such that $f(x^{*}(y),y)=x^{*}(y)$. The following is an assumption on the function $g:\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{2}}$. {assumption} The function $g(x^{*}(\cdot),\cdot):\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{2}}$ is non-expansive, i.e.,
$\|g(x^{*}(y_{1}),y_{1})-g(x^{*}(y_{2}),y_{2})\|\leq\|y_{1}-y_{2}\|,$
for all $y_{1},y_{2}\in\mathbb{R}^{d_{2}}$. Let us denote the set of fixed points of $g(x^{*}(\cdot),\cdot)$ by $\mathcal{Y}^{*}$, i.e., $\mathcal{Y}^{*}=\{y\in\mathbb{R}^{d_{2}}\mid g(x^{*}(y),y)=y\}$. We assume that the set $\mathcal{Y}^{*}$ is non-empty.
Before discussing other required assumptions, we present alternate but equivalent formulations of the above assumptions. Define operator $T_{1}(x,y)=x-f(x,y)$. Then Assumption 2 is equivalent to the operator $T_{1}(x,y)$ being strongly monotone in $x$:
$\langle T_{1}(x_{1},y)-T_{1}(x_{2},y),x_{1}-x_{2}\rangle\geq(1-\mu)\|x_{1}-x_{2}\|^{2}.$
Similarly, define operator $T_{2}(y)=y-g(x^{*}(y),y)$. Then the non-expansiveness in Assumption 2 is equivalent to the operator $T_{2}$ being co-coercive, i.e.,
$\langle T_{2}(y_{1})-T_{2}(y_{2}),y_{1}-y_{2}\rangle\geq\frac{1}{2}\|T_{2}(y_{1})-T_{2}(y_{2})\|^{2}.$
The opposite directions are also true as shown in Lemma D.1 in Appendix D.
Next, we present two assumptions required for our analysis. These are standard assumptions associated with analysis of SA iterations [2, 7]. The first assumption states that the functions $f$ and $g$ are Lipschitz. {assumption} Functions $f(\cdot,\cdot)$ and $g(\cdot,\cdot)$ are $L$-Lipschitz, i.e.,
$\|f(x_{1},y_{1})-f(x_{2},y_{2})\|+\|g(x_{1},y_{1})-g(x_{2},y_{2})\|\leq L(\|x_{1}-x_{2}\|+\|y_{1}-y_{2}\|).$
for all $x_{1},x_{2}\in\mathbb{R}^{d_{1}}$ and $y_{1},y_{2}\in\mathbb{R}^{d_{2}}$. For $x\in\mathbb{R}^{d_{1}}$ and $y\in\mathbb{R}^{d_{2}}$, we also assume that
$\|f(x,y)\|+\|g(x,y)\|\leq\mathfrak{c}_{1}(1+\|x\|+\|y\|).$
The next assumption is that $M_{k+1}$ and $M^{\prime}_{k+1}$ are martingale difference sequences. {assumption} Define the family of $\sigma$-fields $\mathcal{F}_{k}=\sigma(x_{0},y_{0},M_{i},M^{\prime}_{i},i\leq k)$. Then $\{M_{k+1}\}$ and $\{M^{\prime}_{k+1}\}$ are martingale difference sequences with respect to $\mathcal{F}_{k}$, i.e., $\mathbb{E}[M_{k+1}\mid\mathcal{F}_{k}]=\mathbb{E}[M^{\prime}_{k+1}\mid\mathcal{F}_{k}]=0.$ Moreover, $\mathbb{E}[\|M_{k+1}\|^{2}+\|M^{\prime}_{k+1}\|^{2}\mid\mathcal{F}_{k}]\leq\mathfrak{c}_{2}(1+\|x_{k}\|^{2}+\|y_{k}\|^{2}),\;\forall k\geq 0.$
We finally make an assumption about the step size sequences, which is necessary to ensure the time-scale separation between the two updates. {assumption} The step sizes $\alpha_{k}$ and $\beta_{k}$ are of the form
$\alpha_{k}=\frac{\alpha_{0}}{(k+1)^{\mathfrak{a}}}\;\;\text{and}\;\;\beta_{k}=\frac{\beta_{0}}{(k+1)^{\mathfrak{b}}},$
where $10.5<\mathfrak{a}<\mathfrak{b}<$ 1 and $\alpha_{0},\beta_{0}\leq 0.5$. This ensures that they are non-summable $(\sum_{k}\alpha_{k}=\infty,\sum_{k}\beta_{k}=\infty)$, square-summable $(\sum_{k}\alpha^{2}_{k}<\infty,\sum_{k}\beta^{2}_{k}<\infty)$ and non-increasing $(\alpha_{k+1}\leq\alpha_{k}$ and $\beta_{k+1}\leq\beta_{k})$. This also ensures that $\lim_{k\uparrow\infty}\beta_{k}/\alpha_{k}=0$. Finally, we make the key assumption that $\beta_{k}^{2}/\alpha_{k}^{3}\leq 1.$ We restrict our attention to step sizes of this form for simplicity but our results can be extended to other step size sequences. We do not include $1/(k+1)$ in our set of step sizes because it is not optimal for non-expansive fixed point iterations. The assumptions that the step sizes are non-summable, square summable, and satisfy $\beta_{k}/\alpha_{k}$ goes to zero as $k\uparrow\infty$ are standard in stochastic approximation literature [2]. The assumption of $\beta_{k}^{2}/\alpha_{k}^{3}\leq 1$ is the assumption that decides the time-scale separation between the updates and is key for our analysis.
We are now ready to state our main result. An outline of the proof is given in Section 3, and the complete proof is in Appendix A.
Theorem 2.1.
Suppose that Assumptions 2-2 hold. Let $\{x_{k},y_{k}\}$ be generated by (1) with $\beta_{0}/\alpha_{0}\leq(1-\mu)/(2L^{2})$. Then,
  • There exist constants $C_{1},C_{2}>0$ such that for all $k\geq 0$,
    $\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]\leq\frac{C_{1}}{(k+1)^{\mathfrak{a}}}\;\;\text{and}\;\;\mathbb{E}\left[\|g(x^{*}(y_{k}),y_{k})-y_{k}\|^{2}\right]\leq\frac{C_{2}}{(k+1)^{1-\mathfrak{b}}}.$
  • Moreover, $\|x_{k}-x^{*}(y_{k})\|$ converges to zero, and the iterates $y_{k}$ converge to the set $\mathcal{Y}^{*}$ with probability $1$.
Optimal Rate and Step size Choice
The optimal rate for the algorithm based on the above analysis is $1/(k+1)^{0.25-\epsilon}$ where $\epsilon>0$ can be arbitrarily small. This can be achieved by choosing $\mathfrak{a}=0.5+(2/3)\epsilon$ and $\mathfrak{b}=0.75+\epsilon$. Unlike contractive maps, in the analysis of fixed point iterations with non-expansive mappings, a slowly decreasing step size gives a better rate. Hence, a slowly decreasing $\beta_{k}$, or equivalently a smaller $\mathfrak{b}$, gives a better rate. This, along with the assumptions that $\mathfrak{a}>0.5$ and $\beta_{k}^{2}/\alpha_{k}^{3}\leq 1$ (equivalently, $2\mathfrak{b}\geq 3\mathfrak{a}$), dictates the optimal rate.
The assumptions on $\beta_{0}/\alpha_{0}$ are required only for the bounds to hold for all $k\geq 0$. Instead, if we have that $\beta_{k}/\alpha_{k}\leq(1-\mu)/2L^{2}$ holds for all $k\geq K$, then the bounds will hold for all $k\geq K$. The existence of such a $K$ is guaranteed by Assumption 2.
We also want to note that $\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$ and $\mathbb{E}\left[\|g(x^{*}(y_{k}),y_{k})-y_{k}\|^{2}\right]$ are $\mathcal{O}(\alpha_{k})$ and $\mathcal{O}(1/(k\beta_{k}))$, respectively.
2.1 Projected Variant
In this subsection, we consider a variant where the iterate in the faster time-scale is projected to a convex and bounded set at each step.
(2) $x_{k+1}=\mathcal{P}_{\mathcal{X}}\big{(}x_{k}+\alpha_{k}(f(x_{k},y_{k})-x_{k}+M_{k+1})\big{)}$
$y_{k+1}=y_{k}+\beta_{k}(g(x_{k},y_{k})-y_{k}+M^{\prime}_{k+1}).$
Here $\mathcal{X}$ is a convex and bounded set, and $\mathcal{P}_{\mathcal{X}}$ denotes the projection onto set $\mathcal{X}$, i.e., $\mathcal{P}_{\mathcal{X}}(x)=\arg\min_{x^{\prime}\in\mathcal{X}}\|x-x^{\prime}\|$.
We still assume that the function $f(\cdot,y)$ is contractive for all $y$. But the relevant fixed point is now the fixed point for $\mathcal{P}_{\mathcal{X}}(f(\cdot,y))$, i.e., we now define $\hat{x}(y)$ as the unique point that satisfies $\hat{x}(y)=\mathcal{P}_{\mathcal{X}}(f(\hat{x}(y),y))$. In this case, the function $g(\hat{x}(\cdot),\cdot)$ is assumed to be non-expansive. {assumption} The function $g(\hat{x}(\cdot),\cdot):\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{2}}$ is non-expansive, i.e.,
$\|g(\hat{x}(y_{1}),y_{1})-g(\hat{x}(y_{2}),y_{2})\|\leq\|y_{1}-y_{2}\|,$
for all $y_{1},y_{2}\in\mathbb{R}^{d_{2}}$. Let us denote the set of fixed points of $g(\hat{x}(\cdot),\cdot)$ by $\mathcal{\hat{Y}}$, i.e., $\mathcal{\hat{Y}}=\{y\in\mathbb{R}^{d_{2}}\mid g(\hat{x}(y),y)=y\}$. We assume that the set $\mathcal{\hat{Y}}$ is non-empty.
This projected variant is studied as a special case because there exist scenarios where the projection of the faster time-scale updates is why the slower time-scale iteration has a non-expansive mapping (instead of a contractive mapping). Formally, there arise scenarios where the map $g(x^{*}(\cdot),\cdot)$ is contractive, but the map $g(\hat{x}(\cdot),\cdot)$ is not guaranteed to be contractive, and is only non-expansive. We study one such scenario in Subsection 5.3.
The main theorem for this projected variant is similar to Theorem 2.1, and its proof is presented in Appendix B.
Theorem 2.2.
Suppose that Assumptions 2, 2.1 and 2-2 hold. Let $\{x_{k},y_{k}\}$ be generated by (2) with $\beta_{0}/\alpha_{0}\leq(1-\mu)/(2L^{2})$. Then,
  • There exist constants $C_{1}^{\prime},C_{2}^{\prime}>0$ such that for all $k\geq 0$,
    $\mathbb{E}\left[\|x_{k}-\hat{x}(y_{k})\|^{2}\right]\leq\frac{C_{1}^{\prime}}{(k+1)^{\mathfrak{a}}},\;\;\text{and}\;\;\mathbb{E}\left[\|g(\hat{x}(y_{k}),y_{k})-y_{k}\|^{2}\right]\leq\frac{C_{2}^{\prime}}{(k+1)^{1-\mathfrak{b}}}.$
  • Moreover, $\|x_{k}-\hat{x}(y_{k})\|$ converges to zero, and the iterates $y_{k}$ converge to the set $\mathcal{\hat{Y}}$ with probability $1$.
3 Proof Outline
In this section, we give a rough sketch for the proof for Theorem 2.1. Appendix A presents the complete proof. The proof takes inspiration from analysis of two-time-scale iterations with contractive maps [10] and analysis of inexact Krasnoselskii-Mann (KM) iterations [9, 19], but requires novel analysis to deal with the additional errors of two-time-scale updates in a stochastic setting with non-expansive maps. For simplicity, we define $h(y)=g(x^{*}(y),y)$.
The first step in the proof is to show that $x^{*}(y)$ is Lipschitz in $y$.
Lemma 3.1.
Suppose Assumptions 2 and 2 hold. Then $x^{*}(y)$, the unique fixed point of $f(\cdot,y)$, satisfies
$\|x^{*}(y_{1})-x^{*}(y_{2})\|\leq L_{0}\|y_{1}-y_{2}\|,$
for all $y_{1},y_{2}\in\mathbb{R}^{d_{2}}$. Here $L_{0}\coloneqq L/(1-\mu)$ is the Lipschitz parameter for the mapping $x^{*}(\cdot):\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{1}}$.
This lemma ensures that the target points for the faster iteration $x^{*}(y_{k})$ move slowly as the iteration $\{y_{k}\}$ moves slowly. Apart from the noise sequence $M_{k+1}$, ($x^{*}(y_{k+1})-x^{*}(y_{k})$) acts as another ‘error term’ for the faster iteration. The Lipschitz nature of $x^{*}(\cdot)$, along with the slow changes in $y_{k}$, guarantees that this error term is small.
The next step of the proof is to obtain an intermediate bound for $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}]$ and $\mathbb{E}[\|y_{k}-y^{*}\|^{2}]$. Here $y^{*}\in\mathcal{Y}^{*}$ is some fixed point for the mapping $g(x^{*}(\cdot),\cdot)$.
Lemma 3.2.
Suppose the setting of Theorem 2.1 holds. Then for $y^{*}\in\mathcal{Y}^{*}$, iterates $\{x_{k},y_{k}\}$ satisfy the following bounds.
$\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]$ $\leq(1-\mu^{\prime}\alpha_{k})\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$
$\;\;+\Gamma_{1}\alpha_{k}^{2}\left(1+\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]+\mathbb{E}\left[\|y_{k}-y^{*}\|^{2}\right]\right),$
$\mathbb{E}\left[\|y_{k+1}-y^{*}\|^{2}\right]$ $\leq\mathbb{E}\left[\|y_{k}-y^{*}\|^{2}\right]+\mu^{\prime}\alpha_{k}\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$
$\;\;+\Gamma_{2}\alpha_{k}^{2}\left(1+\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]+\mathbb{E}\left[\|y_{k}-y^{*}\|^{2}\right]\right)$
$\;\;-\beta_{k}(1-\beta_{k})\mathbb{E}\left[\|y_{k}-g(x_{k},y_{k})\|^{2}\right].$
Here $\Gamma_{1},\Gamma_{2}$ are some constants and $\mu^{\prime}=1-\mu$.
The assumption that $\beta_{k}^{2}\leq\alpha_{k}^{3}$ is key for obtaining the above intermediate bounds. We first add the two intermediate bounds. Then we show that the iterates are bounded in expectation using the assumption that the step size sequence $\alpha_{k}$ is square-summable.
Lemma 3.3.
Suppose the setting of Theorem 2.1 holds. Then the iterates $x_{k}$ and $y_{k}$ are bounded in expectation. Specifically, there exists $\Gamma_{3},\kappa_{1}>0$ such that for all $k\geq 0$, and for $y^{*}\in\mathcal{Y}^{*}$,
$\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2}\right]\leq\Gamma_{3}\coloneqq\kappa_{1}\left(\kappa_{1}+\|x_{0}-x^{*}(y_{0})\|^{2}+\|y_{0}-y^{*}\|^{2}\right).$
The above lemmas are sufficient to prove that iterates converge to the set of fixed points w.p. 1. For this, we use the ODE approach for analysis of SA algorithms. A key requirement for the approach is almost sure finiteness of iterates, which follows from Lemma 3.3. The bound for $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}]$ is also direct from the above lemma. We get the following recursive relation using Lemmas 3.2 and 3.3.
$\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]\leq(1-\mu^{\prime}\alpha_{k})\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]+\Gamma_{1}(1+\Gamma_{3})\alpha_{k}^{2}.$
Solving this recursive relation gives us the bound that $\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]=\mathcal{O}(\alpha_{k})$. Similarly, on analyzing the other bound, we get that the summation of $\beta_{i}\mathbb{E}\left[\|y_{i}-h(y_{i})\|^{2}\right]$ over all $i$ is bounded by a constant. Recall that $h(y)=g(x^{*}(y),y)$.
Lemma 3.4.
Suppose the setting of Theorem 2.1 holds. Then,
  1. a)
    $\|x_{k}-x^{*}(y_{k})\|$ converges to zero, and $y_{k}$ converges to the set $\mathcal{Y}^{*}$ w.p. 1.
  2. b)
    There exist constants $C_{1},\Gamma_{4}>0$ such that, for all $k\geq 0$,
    $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}]\leq\frac{C_{1}}{(k+1)^{\mathfrak{a}}}\;\;\text{and}\;\;\sum_{i=0}^{k}\beta_{i}\mathbb{E}\left[\|h(y_{i})-y_{i}\|^{2}\right]\leq\Gamma_{4},$
    where $C_{1}=\kappa_{2}\left(1+\|x_{0}-x^{*}(y_{0})\|^{2}+\|y_{0}-y^{*}\|^{2}\right)$ for some $\kappa_{2}>0$.
The above lemma would be sufficient if we wanted to get a bound on the term $\min_{0\leq i\leq k}\mathbb{E}\left[\|h(y_{i})-y_{i}\|^{2}\right]$. But we wish to get a bound on $\mathbb{E}\left[\|h(y_{k})-y_{k}\|^{2}\right]$.
For this, we would like to get a bound on $\mathbb{E}[\|h(y_{i})-y_{i}\|^{2}]$ in terms of $\mathbb{E}[\|h(y_{k})-y_{k}\|^{2}]$. But the martingale noise $M^{\prime}_{k+1}$ does not allow us to obtain a useful bound. To solve this, we define an averaged error sequence $U_{k+1}=(1-\beta_{k})U_{k}+\beta_{k}M^{\prime}_{k+1}$, with $U_{0}=0$, and a modified iteration of $z_{k}=y_{k}-U_{k}$. We first show that $\mathbb{E}[\|y_{k}-z_{k}\|^{2}]$ decays at a sufficiently fast rate, implying that we can analyze $\mathbb{E}[h(z_{k})-z_{k}\|^{2}]$ instead of $\mathbb{E}[\|h(y_{k})-y_{k}\|^{2}].$ Next, we study the iteration for $z_{k}$, showing that the error sequence for $z_{k}$ satisfies a similar boundedness condition for the summation as in Lemma 3.4 b). Finally, we fulfill the purpose of studying $z_{k}$ instead of $y_{k}$, by obtaining a bound on $\mathbb{E}[\|h(z_{i})-z_{i}\|^{2}]$ in terms of $\mathbb{E}[\|h(z_{k})-z_{k}\|^{2}]$.
Lemma 3.5.
Suppose the setting of Theorem 2.6 holds. Then, for all $k\geq 0$,
  1. a)
    $\mathbb{E}[\|y_{k}-z_{k}\|^{2}]=\mathbb{E}[\|U_{k}\|^{2}]\leq\Gamma_{5}\beta_{k}$, for some $\Gamma_{5}>0$.
  2. b)
    $\sum_{i=0}^{k}\beta_{i}\mathbb{E}\left[\|h(z_{i})-z_{i}\|^{2}\right]\leq\Gamma_{6}$, for some $\Gamma_{6}>0$.
  3. c)
    For some $\Gamma_{7}>0$,
    $\mathbb{E}\left[\|h(z_{i})-z_{i}\|^{2}\right]\geq\mathbb{E}\left[\|h(z_{k})-z_{k}\|^{2}\right]-\sum_{j=i}^{k-1}\Gamma_{7}\beta_{j}\alpha_{j}.$
Lemma 3.5 implies that
$\mathbb{E}\left[\|h(z_{k})-z_{k}\|^{2}\right]\sum_{i=0}^{k}\beta_{i}\leq\Gamma_{6}+\Gamma_{7}\sum_{i=0}^{k}\beta_{i}\sum_{j=i}^{k-1}\beta_{j}\alpha_{j}.$
We finally show that the summation on the right-hand side above is bounded. This requires the assumptions of square-summability of $\alpha_{k}$ and $\beta_{k}^{2}\leq\alpha_{k}^{3}$ and completes the proof for Theorem 2.1.
4 Gradient Descent Operators in Slower Time-Scale
An important example of the iteration in (1) is one where the slower time-scale (iteration for $y_{k}$) is a gradient descent iteration in the limit $x_{k}\rightarrow x^{*}(y_{k})$. The following assumption formally states this. {assumption} We assume that there exists differentiable function $J:\mathbb{R}^{d_{2}}\mapsto\mathbb{R}$ such that $g(x^{*}(y),y)-y=-\nabla J(y)$. We further assume that $J(\cdot)$ is bounded from below, i.e., there exists $J_{min}\in\mathbb{R}$ such that $J(y)\geq J_{min}$ for all $y$. Depending on the assumptions on function $J(\cdot)$, different types of convergence rates can be obtained. The function $J(\cdot)$ is smooth in all cases (i.e., $\nabla J(\cdot)$ is Lipschitz).
$J(\cdot)$ is strongly convex
In this case, the function $g(x^{*}(\cdot),\cdot)$ can be shown to be contractive, and hence we have a two-time-scale iteration where both time-scales have contractive mappings. In this case, the result from [10] can directly be applied, to obtain a bound of $\mathcal{O}(1/k^{2/3})$ on $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2}]$. Here $y^{*}$ is the unique minimizer for $J(\cdot)$.
$J(\cdot)$ is convex
In this case, the function $g(x^{*}(\cdot),\cdot)$ is non-expansive. This gives us a two-time-scale iteration, which satisfies the assumptions for Theorem 2.1. We can apply our result to get a bound of $\mathcal{O}(1/k^{0.25-\epsilon})$ on $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}+\|\nabla J(y_{k})\|^{2}]$. This case is discussed in more detail in Subsection 5.1, when we discuss minimax optimization.
$J(\cdot)$ is non-convex
In this case, the function $g(x^{*}(\cdot),\cdot)$ is not necessarily non-expansive, and hence Theorem 2.1 can not be applied. But under Lipschitz assumptions on function $g(\cdot,\cdot)$, the function $J(\cdot)$ is smooth, and hence bounds can be obtained in this case using a proof technique similar to the proof for Theorem 2.1. Consider the following assumption. {assumption} We assume that the following statements hold.
  1. a)
    Function $f(\cdot,y)$ is contractive with parameter $\mu$ (Assumption 2).
  2. b)
    Function $f$ and $g$ are Lipschitz with parameter $L$ (Assumption 2).
  3. c)
    $M_{k+1}$ and $M^{\prime}_{k+1}$ are martingale difference sequences with bounded variances, i.e., $\mathbb{E}[\|M_{k+1}\|^{2}+\|M^{\prime}_{k+1}\|^{2}\mid\mathcal{F}_{k}]\leq\mathfrak{c}_{2}$.
  4. d)
    The stepsize sequences satisfy $\mathfrak{a}<\mathfrak{b}<1$ and $0.5<\mathfrak{b}$. We also have the key assumption that $\beta_{k}^{2}\leq\alpha_{k}^{3}$ and $\beta_{k}\alpha_{k}\leq\beta_{0}\alpha_{0}/(k+1)$.
The key difference in assumptions is that we no longer require $\alpha_{k}$ to be square summable, allowing sequences such as $\alpha_{k}=\mathcal{O}(1/k^{0.4})$ and $\beta_{k}=\mathcal{O}(1/k^{0.6})$. Then, we have the following theorem, with its proof presented in Appendix C.
Theorem 4.1.
Suppose that Assumptions 4 and 4 hold. Then there exist constants $C_{3},C_{4}$ such that if $\alpha_{0}\leq C_{4}$, then for all $k\geq 0$,
$\min_{0\leq i\leq k}\mathbb{E}\left[\|\nabla J(y_{i})\|^{2}\right]\leq\frac{C_{3}(1+\log(k+1))}{(1+k)^{1-\mathfrak{b}}}.$
The optimal rate in this case is obtained at $\mathfrak{a}=0.4$ and $\mathfrak{b}=0.6$, with a bound of $\min_{0\leq i\leq k}\mathbb{E}\left[\|\nabla J(y_{i})\|^{2}\right]=\mathcal{O}(\log(k)/k^{0.4})$. In this case, we do not have a bound on the gradient at time $k$, but only have a bound on the minimum gradient attained before time $k$. Such non-convex two-time-scale algorithms arise in non-convex minimax optimization and reinforcement learning (actor-critic algorithms) with similar bounds obtained in [29, 28] under different settings.
5 Applications
In this section, we discuss three applications of the framework and results discussed in Section 2.
5.1 Minimax Optimization
Consider the following minimax optimization problem:
$\min_{y\in\mathbb{R}^{d_{2}}}\max_{x\in\mathbb{R}^{d_{1}}}H(x,y),$
where $H:\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\mapsto\mathbb{R}$ is the objective function. This formulation of finding saddle point solutions to minimax optimization problems arises in several areas, including two-player zero-sum games, generative adversarial networks, and robust statistics [20]. We consider the Two-Time-scale Stochastic Gradient Descent Ascent (TTSGDA) algorithm [20] for finding saddle points.
(3) $x_{k+1}=x_{k}+\alpha_{k}(\nabla_{x}H(x_{k},y_{k})+M_{k+1})$
$y_{k+1}=y_{k}+\beta_{k}(-\nabla_{y}H(x_{k},y_{k})+M^{\prime}_{k+1}).$
Here $M_{k+1}$ and $M^{\prime}_{k+1}$ denote the noise arising from taking gradient samples at the current iterate. These noise sequences are assumed to be unbiased and have bounded variance, i.e., they satisfy Assumption 2 [20].
We work under the setting of smooth strongly concave-convex minimax optimization. This is formally stated in the following assumption. {assumption} Function $H(x,y)$ is $L-$smooth, $\rho-$strongly concave in $x$ and convex in $y$. This assumption is sufficient to ensure that the faster time-scale (update of iterate $x_{k}$) has a contractive mapping and the slower time-scale (update of iterate $y_{k}$) has a non-expansive mapping. Define $\Phi(y)=\max_{x\in\mathbb{R}^{d_{1}}}H(x,y)$. Then we have the following result, with its proof presented in Appendix D.
Corollary 5.1.
Suppose Assumption 5.1 holds for iterates $\{x_{k},y_{k}\}$ defined in (3), and let $\alpha_{k}=1/(k+1)^{0.5+0.66\epsilon}$ and $\beta_{k}=1/(k+1)^{0.75+\epsilon}$. Then the iterates $\{x_{k},y_{k}\}$ converge to the set of saddle points of the function $H(x,y)$ a.s., and $\mathbb{E}\left[\|\nabla\Phi(y)\|^{2}\right]\leq\mathcal{O}(1/k^{1/4-\epsilon}).$ Here $\epsilon$ can be arbitrarily small.
We also want to mention that a similar result can be obtained for the case where $x_{k}$ is bounded in a convex and bounded set $\mathcal{X}$ using the framework from Subsection 2.1.
5.2 Linear Stochastic Approximation
Consider the problem of finding solutions for the following set of linear equations:
(4) $A_{11}x+A_{12}y=b_{1}\quad\text{and}\quad A_{21}x+A_{22}y=b_{2}.$
At each time $k$, we obtain unbiased estimates $\tilde{A}_{ij}^{(k+1)}$ and $\tilde{b}_{i}^{(k+1)}$ of each $A_{ij}$ and $b_{i}$, $i,j\in\{1,2\}$. This problem is solved using a linear two-time-scale algorithm.
(5) $x_{k+1}=x_{k}+\alpha_{k}\left(\tilde{b}_{1}^{(k+1)}-\tilde{A}_{11}^{(k+1)}x_{k}-\tilde{A}_{12}^{(k+1)}y_{k}\right)$
$y_{k+1}=y_{k}+\beta_{k}\left(\tilde{b}_{2}^{(k+1)}-\tilde{A}_{21}^{(k+1)}x_{k}-\tilde{A}_{22}^{(k+1)}y_{k}\right).$
This algorithm finds use in reinforcement learning [16] and linear SA with Polyak averaging [18]. Define matrix $\Delta=A_{22}-A_{21}A_{11}^{-1}A_{12}$. Under the assumption that matrices $-A_{11}$ and $-\Delta$ are Hurwitz, i.e., the real part of all their eigenvalues is negative, iteration (5) can be written as two-time-scale SA with both time-scales as fixed point iterations with contractive mappings.
In this paper, we consider the case where the matrix $-A_{11}$ is negative definite and the matrix $-\Delta$ is negative semidefinite. Formally, we make the following assumption. {assumption} Iteration (5) satisfies the following.
  1. a)
    The matrix $-A_{11}$ is negative definite and the matrix $-\Delta$ is (non-zero) negative semidefinite.
  2. b)
    There exists a solution for the system of equations in (4).
  3. c)
    For $i,j\in\{1,2\}$, $\mathbb{E}\left[\tilde{A}_{ij}^{(k+1)}\mid\mathcal{F}_{k}\right]=A_{ij},\mathbb{E}\left[\tilde{b}_{i}^{(k+1)}\mid\mathcal{F}_{k}\right]=b_{i}.$ Moreover,
    $\mathbb{E}\left[\|\tilde{A}_{ij}^{(k+1)}-A_{ij}\|^{2}\mid\mathcal{F}_{k}\right]\leq\mathfrak{c}_{3},\quad\text{and}\quad\mathbb{E}\left[\|\tilde{b}_{i}^{(k+1)}-b_{i}\|^{2}\mid\mathcal{F}_{k}\right]\leq\mathfrak{c}_{3}.$
Part a) of this assumption ensures that iteration (5) satisfies Assumptions 2 and 2. As Hurwitz matrices also correspond to contractive fixed point iterations [18], we can also work with Hurwitz $-A_{11}$ (instead of negative definite). Here is the result for this setting, with its proof presented in Appendix D.
Corollary 5.2.
Suppose Assumption 5.2 holds for iterates $\{x_{k},y_{k}\}$ defined in (5), and let $\alpha_{k}=1/(k+1)^{0.5+0.66\epsilon}$ and $\beta_{k}=1/(k+1)^{0.75+\epsilon}$. Then the iterates $\{x_{k},y_{k}\}$ converge to the set of solutions of system of equations (4) a.s., and
$\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]\leq\mathcal{O}\left(\frac{1}{k^{0.5}}\right)\;\;\text{and}\;\;\mathbb{E}\left[\|A_{21}x_{k}+A_{22}y_{k}-b_{2}\|^{2}\right]\leq\mathcal{O}\left(\frac{1}{k^{0.25-\epsilon}}\right).$
Here $\epsilon$ can be arbitrarily small, and $x^{*}(y)=A_{11}^{-1}(b_{1}-A_{12}y)$.
5.3 Constrained Optimization with Lagrangian Multipliers
Consider the following constrained optimization problem for $x\in\mathbb{R}^{d_{1}}$:
$\text{maximize}\;\;H_{0}(x)$
subject to:
$\;\;\;\;\;H_{i}(x)\leq 0,i=1,\ldots,m$
(6) $\;\;\;\;\;Ax=b_{0}.$
Here the constraints $H_{i}(x)\leq 0$ together represent the set $\mathcal{X}$, i.e., $\mathcal{X}=\{x\mid H_{i}(x)\leq 0,i=1,\ldots,m\}.$ The equation $Ax=b_{0}$ for $A\in\mathbb{R}^{d_{2}\times d_{1}}$ and $b_{0}\in\mathbb{R}^{d_{2}}$ represents the additional linear constraint. This problem can be solved using Lagrange multipliers, specifically, a two-time-scale Lagrangian optimization method [23].
Consider the following iteration.
(7) $x_{k+1}=\mathcal{P}_{\mathcal{X}}\big{(}x_{k}+\alpha_{k}(\nabla H_{0}(x_{k})-A^{T}\lambda_{k}+M_{k+1})\big{)}$
$\lambda_{k+1}=\lambda_{k}+\beta_{k}(Ax_{k}-b_{0}).$
Here $\lambda_{k}\in\mathbb{R}^{d_{2}}$ denotes the Lagrange multiplier. $M_{k+1}$ denotes the noise arising from taking gradient samples at the current value of $x_{k}$, and satisfies Assumption 2.
This algorithm is particularly useful in distributed settings, where each node or agent updates some part of the variable $x$ under some local convex constraints, and there is a global linear constraint. In that case, the local updates only require knowledge of the gradient, the Lagrange multiplier, and their local constraints. Another similar setting in which a similar algorithm can be used is generalized Nash equilibrium problems (or Nash equilibrium under constraints) [14] and game control for strongly monotone games [5]. In this setting, each player takes action in a convex set and updates their action using gradient ascent. The game controller has a global linear objective which they want the player’s actions to satisfy [5].
We make the following assumption. The second part of the assumption is required for the existence of $\lambda$ which satisfies $A\hat{x}(\lambda)=b_{0}$. {assumption} Iteration (7) satisfies the following statements.
  1. a)
    Function $H_{0}(x)$ is $\rho-$strongly concave and $L$-smooth.
  2. b)
    Functions $H_{i}(x)$ are continuously differentiable and convex. Moreover, there exists $x\in\text{int}(\mathcal{X})$ (i.e., $H_{i}(x)<0,i=1,\ldots,m$) such that $Ax=b_{0}$.
This problem is an example of a setting where the projection in the faster time-scale gives rise to the non-expansiveness in the slower time-scale. Recall the definitions of $x^{*}(y)$ and $\hat{x}(y)$. In this case, $x^{*}(\lambda)$ is the unique solution to $\nabla H_{0}(x)-A^{T}\lambda=0$ and $\hat{x}(\lambda)$ is the unique solution to $x=\mathcal{P}_{\mathcal{X}}(x+\nabla H_{0}(x_{k})-A^{T}\lambda)$. Then we have the following lemma.
Lemma 5.3.
Under Assumption 5.3 and the assumption that the matrix $A$ has full row rank:
  1. a)
    $-Ax^{*}(\lambda)$ is strongly monotone.
  2. b)
    $-A\hat{x}(\lambda)$ is co-coercive.
The above lemma implies that in the absence of the projection onto set $\mathcal{X}$, iteration (7) would have contractive mappings in both time-scales. On the other hand, the slower time-scale has a non-expansive mapping in the presence of the projection. The following is our main result for this setting.
Corollary 5.4.
Suppose Assumption 5.3 holds for iterates $\{x_{k},\lambda_{k}\}$ defined in (7), and let $\alpha_{k}=1/(k+1)^{0.5+0.66\epsilon}$ and $\beta_{k}=1/(k+1)^{0.75+\epsilon}$. Then $x_{k}$ converges to $\hat{x}(\lambda_{k})$ and $Ax_{k}$ converges to $b_{0}$ a.s., and
$\mathbb{E}\left[\|x_{k}-\hat{x}(y_{k})\|^{2}\right]\leq\mathcal{O}\left(\frac{1}{k^{0.5}}\right)\;\;\text{and}\;\;\mathbb{E}\left[\|Ax_{k}-b_{0}\|^{2}\right]\leq\mathcal{O}\left(\frac{1}{k^{0.25-\epsilon}}\right).$
Here $\epsilon$ can be arbitrarily small.
6 Conclusion
In this paper, we studied two-time-scale iterations where the slower time-scale has a non-expansive mapping. Our main contribution is the finite-time bound for this setting. We showed that the mean square error decays at a rate of $\mathcal{O}(k^{-1/4+\epsilon})$, where $\epsilon>0$ is arbitrarily small depending on the step size. Our analysis takes inspiration from techniques in contractive SA and inexact KM iterations. Extensions of our work include studying the same algorithm under Markovian noise and extending the result to Banach spaces. Another extension could be to study non-expansive mappings in both time-scales, but the analysis would require further assumptions to ensure that the faster time-scale has a unique fixed point for each $y$.
Appendix A Proofs required for Theorem 2.1
A.1 Proof for Lemma 3.1
Proof A.1 (Proof for Lemma 3.1).
Note that
$\|x^{*}(y_{1})-x^{*}(y_{2})\|$ $=\|f(x^{*}(y_{1}),y_{1})-f(x^{*}(y_{2}),y_{2})\|$
$\leq\|f(x^{*}(y_{1}),y_{1})-f(x^{*}(y_{2}),y_{1})\|+\|f(x^{*}(y_{2}),y_{1})-f(x^{*}(y_{2}),y_{2})\|$
$\leq\mu\|x^{*}(y_{1})-x^{*}(y_{2})\|+L\|y_{1}-y_{2}\|.$
Here the first equality follows from the fact that $x^{*}(y)$ is a fixed point for $f(\cdot,y)$ and the last inequality follows from contractiveness and Lipschitzness of $f(x,y)$ in $x$ and $y$, respectively. This implies that $(1-\mu)\|x^{*}(y_{1})-x^{*}(y_{2})\|\leq L\|y_{1}-y_{2}\|.$ Hence $x^{*}(\cdot)$ is Lipschitz with constant $L_{0}\coloneqq L/\mu^{\prime}$, where $\mu^{\prime}=1-\mu$.
A.2 Proof for Lemma 3.2
Proof A.2 (Proof for Lemma 3.2).
For this proof, we first give an intermediate bound on $\mathbb{E}[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}]$ and then on $\mathbb{E}[\|y_{k+1}-y^{*}\|^{2}]$.
Recursive Bound on $\mathbb{E}[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}]$
By adding and subtracting $x^{*}(y_{k})$,
$\|x_{k+1}-x^{*}(y_{k+1})\|^{2}$
(8a) $=\|x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k})\|^{2}$
(8b) $\;\;+\|\alpha_{k}M_{k+1}+x^{*}(y_{k})-x^{*}(y_{k+1})\|^{2}$
(8c) $\;\;+2\Big{\langle}x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k}),\alpha_{k}M_{k+1}+x^{*}(y_{k})-x^{*}(y_{k+1})\Big{\rangle}.$
We deal with the three terms (8a)-(8c) separately as follows.
Term 8a
$\|x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k})\|^{2}$
$=\|(1-\alpha_{k})(x_{k}-x^{*}(y_{k}))+\alpha_{k}(f(x_{k},y_{k})-f(x^{*}(y_{k}),y_{k}))\|^{2}$
$\leq\big{(}(1-\alpha_{k})\|x_{k}-x^{*}(y_{k})\|+\alpha_{k}\mu\|x_{k}-x^{*}(y_{k})\|\big{)}^{2}$
(9) $\leq(1-\mu^{\prime}\alpha_{k})^{2}\|x_{k}-x^{*}(y_{k})\|^{2}.$
Here, the first equality follows from the fact that $f(x^{*}(y_{k}),y_{k})=x^{*}(y_{k})$, and the first inequality follows from the property that $f(\cdot,y)$ is contractive.
Term 8b This term can be thought of as the ‘error term’ for the iteration with the faster time-scale. We simplify it as follows.
$\mathbb{E}[\|\alpha_{k}M_{k+1}+x^{*}(y_{k})-x^{*}(y_{k+1})\|^{2}\mid\mathcal{F}_{k}]$
$\leq\mathbb{E}\left[2\alpha_{k}^{2}\|M_{k+1}\|^{2}+2\|x^{*}(y_{k+1})-x^{*}(y_{k})\|^{2}\mid\mathcal{F}_{k}\right]$
$\leq 2\mathfrak{c}_{2}\alpha_{k}^{2}(1+\|x_{k}\|^{2}+\|y_{k}\|^{2})+2L_{0}^{2}\zeta_{1}\beta_{k}^{2}(1+\|x_{k}\|^{2}+\|y_{k}\|^{2})$
(10) $=(2\mathfrak{c}_{2}\alpha_{k}^{2}+2L_{0}^{2}\zeta_{1}\beta_{k}^{2})(1+\|x_{k}\|^{2}+\|y_{k}\|^{2}),$
where the second inequality uses Assumption 2 and Lemma A.7a) (Appendix A.7).
Term 8c To simplify (8c), we first note that:
$2\Big{\langle}x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k}),x^{*}(y_{k})-x^{*}(y_{k+1})\Big{\rangle}$
$\leq\mu^{\prime}\alpha_{k}\|x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k})\|^{2}+\frac{1}{\mu^{\prime}\alpha_{k}}\|x^{*}(y_{k})-x^{*}(y_{k+1})\|^{2}.$
This follows from applying Cauchy-Schwarz inequality, and then applying AM-GM inequality - $2ab\leq\vartheta a^{2}+(1/\vartheta)b^{2}$, with $\vartheta=\mu^{\prime}\alpha_{k}$. We apply (A.2) and Lemma A.7a), respectively, for the two terms. Using the fact that $M_{k+1}$ is a martingale difference sequence, we have $\mathbb{E}\Big{[}2\big{\langle}x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k}),M_{k+1}\big{\rangle}\mid\mathcal{F}_{k}\Big{]}=0.$ Hence,
$\mathbb{E}\Big{[}2\big{\langle}x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k}),\alpha_{k}M_{k+1}+x^{*}(y_{k})-x^{*}(y_{k+1})\big{\rangle}\mid\mathcal{F}_{k}\Big{]}$
(11) $\leq\mu^{\prime}\alpha_{k}\|x_{k}-x^{*}(y_{k})\|^{2}+\frac{L_{0}^{2}\zeta_{1}}{\mu^{\prime}}\frac{\beta_{k}^{2}}{\alpha_{k}}(1+\|x_{k}\|^{2}+\|y_{k}\|^{2}).$
Having obtained a bound on the conditional expectation of the three terms in (8), we now combine the bounds (A.2)-(A.2) after taking expectation to get
$\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]$ $\leq\left((1-\mu^{\prime}\alpha_{k})^{2}+\mu^{\prime}\alpha_{k}\right)\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$
$\;\;+\left(2\mathfrak{c}_{2}\alpha_{k}^{2}+2L_{0}^{2}\zeta_{1}\beta_{k}^{2}+\frac{L_{0}^{2}\zeta_{1}}{\mu^{\prime}}\frac{\beta_{k}^{2}}{\alpha_{k}}\right)\mathbb{E}\left[1+\|x_{k}\|^{2}+\|y_{k}\|^{2}\right].$
Now for $y^{*}\in\mathcal{Y}^{*}$, we apply Lemma A.7b).
$\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]$
$\leq\left(1-2\mu^{\prime}\alpha_{k}+\mu^{\prime 2}\alpha_{k}^{2}+\mu^{\prime}\alpha_{k}\right)\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$
$\;\;+\left(2\mathfrak{c}_{2}\alpha_{k}^{2}+2L_{0}^{2}\zeta_{1}\beta_{k}^{2}+\frac{L_{0}^{2}\zeta_{1}}{\mu^{\prime}}\frac{\beta_{k}^{2}}{\alpha_{k}}\right)\zeta_{2}\left(1+\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]+\mathbb{E}\left[\|y_{k}-y^{*}\|^{2}\right]\right).$
Now, we use our assumption that $\beta_{k}^{2}/\alpha_{k}^{3}\leq 1$ (Assumption 2) to obtain
$\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]$ $\leq(1-\mu^{\prime}\alpha_{k})\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$
(12) $\;\;+\Gamma_{1}\alpha_{k}^{2}\left(1+\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]+\mathbb{E}\left[\|y_{k}-y^{*}\|^{2}\right]\right),$
where $\Gamma_{1}\coloneqq\left(\mu^{\prime 2}+2\mathfrak{c}_{2}\zeta_{2}+2L_{0}^{2}\zeta_{1}\zeta_{2}+2L_{0}^{2}\zeta_{1}\zeta_{2}/\mu^{\prime}\right)$. This is our intermediate recursive bound on $\mathbb{E}[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}]$.
Recursive Bound on $\mathbb{E}[\|y_{k+1}-y^{*}\|^{2}]$
We first use Corollary 2.14 from [1] which states that for any $\epsilon\in(0,1)$ and $y_{1},y_{2}\in\mathbb{R}^{d_{2}}$, we have $\|(1-\epsilon)y_{1}+\epsilon y_{2}\|^{2}=(1-\epsilon)\|y_{1}\|^{2}+\epsilon\|y_{2}\|^{2}-\epsilon(1-\epsilon)\|y_{1}-y_{2}\|^{2}.$
$\|y_{k+1}-y^{*}\|^{2}$ $=\|(1-\beta_{k})(y_{k}-y^{*})+\beta_{k}(g(x_{k},y_{k})-y^{*}+M^{\prime}_{k+1})\|^{2}$
(13a) $=(1-\beta_{k})\|y_{k}-y^{*}\|^{2}+\beta_{k}\|g(x_{k},y_{k})-y^{*}+M^{\prime}_{k+1}\|^{2}$
(13b) $\;\;-(1-\beta_{k})(\beta_{k})\|y_{k}-g(x_{k},y_{k})-M^{\prime}_{k+1}\|^{2}.$
Term 13a For the term (13a), we first note that
$\|g(x_{k},y_{k})-y^{*}+M^{\prime}_{k+1}\|^{2}$ $=\|g(x_{k},y_{k})-y^{*}\|^{2}+\|M^{\prime}_{k+1}\|^{2}+2\langle g(x_{k},y_{k})-y^{*},M^{\prime}_{k+1}\rangle.$
The expectation of the third term above is zero when conditioned on $\mathcal{F}_{k}$. Now, recall that $h(y)=g(x^{*}(y),y)$ is non-expansive. For the first term, we note the following.
$\|g(x_{k},y_{k})-y^{*}\|^{2}=\|h(y_{k})-y^{*}+g(x_{k},y_{k})-h(y_{k})\|^{2}$
$=\|h(y_{k})-y^{*}\|^{2}+\|g(x_{k},y_{k})-h(y_{k})\|^{2}+2\langle h(y_{k})-y^{*},g(x_{k},y_{k})-h(y_{k})\rangle$
$\leq\|h(y_{k})-h(y^{*})\|^{2}+\|g(x_{k},y_{k})-h(y_{k})\|^{2}+2\|h(y_{k})-h(y^{*})\|\|g(x_{k},y_{k})-h(y_{k})\|$
$\leq\|y_{k}-y^{*}\|^{2}+L^{2}\|x_{k}-x^{*}(y_{k})\|^{2}+2L\|y_{k}-y^{*}\|\|x_{k}-x^{*}(y_{k})\|.$
Here, the last inequality follows from the non-expansiveness of map $h(\cdot)$ (Assumption 2), and the Lipschitz nature of map $g(\cdot,\cdot)$ (Assumption 2). For the last term above, we use Young’s inequality to note the following.
$2L\beta_{k}\|y_{k}-y^{*}\|\|x_{k}-x^{*}(y_{k})\|\leq\frac{\mu^{\prime}\alpha_{k}}{2}\|x_{k}-x^{*}(y_{k})\|^{2}+\frac{2L^{2}\beta_{k}^{2}}{\mu^{\prime}\alpha_{k}}\|y_{k}-y^{*}\|^{2}.$
Combining these results, we finally get the following result.
$(1-\beta_{k})\mathbb{E}[\|y_{k}-y^{*}\|^{2}]+\beta_{k}\mathbb{E}[\|g(x_{k},y_{k})-y^{*}+M^{\prime}_{k+1}\|^{2}\mid\mathcal{F}_{k}]$
$\leq\left(L^{2}\beta_{k}+\frac{\mu^{\prime}\alpha_{k}}{2}\right)\|x_{k}-x^{*}(y_{k})\|^{2}+\left(1+\frac{2L^{2}\beta_{k}^{2}}{\mu^{\prime}\alpha_{k}}\right)\|y_{k}-y^{*}\|^{2}+\beta_{k}\mathbb{E}[\|M^{\prime}_{k+1}\|^{2}\mid\mathcal{F}_{k}].$
Term 13b We first note that
$\|y_{k}-g(x_{k},y_{k})-M^{\prime}_{k+1}\|^{2}=\|y_{k}-g(x_{k},y_{k})\|^{2}+\|M^{\prime}_{k+1}\|^{2}-2\langle y_{k}-g(x_{k},y_{k}),M^{\prime}_{k+1}\rangle.$
As $M^{\prime}_{k+1}$ is a martingale difference sequence, the expectation of the third term is zero.
$-\beta_{k}(1-\beta_{k})\mathbb{E}[\|y_{k}-g(x_{k},y_{k})-M^{\prime}_{k+1}\|^{2}\mid\mathcal{F}_{k}]$
$=-\beta_{k}(1-\beta_{k})\|y_{k}-g(x_{k},y_{k})\|^{2}+\left(-\beta_{k}+\beta_{k}^{2}\right)\mathbb{E}[\|M^{\prime}_{k+1}\|^{2}\mid\mathcal{F}_{k}].$
Combining the bounds for terms (13a) and (13b), we now have
$\mathbb{E}[\|y_{k+1}-y^{*}\|^{2}\mid\mathcal{F}_{k}]$ $\leq-\beta_{k}(1-\beta_{k})\|y_{k}-g(x_{k},y_{k})\|^{2}+\beta_{k}^{2}\mathbb{E}[\|M^{\prime}_{k+1}\|^{2}\mid\mathcal{F}_{k}]$
$\;\;+\left(1+\frac{2L^{2}\beta_{k}^{2}}{\mu^{\prime}\alpha_{k}}\right)\|y_{k}-y^{*}\|^{2}+\left(L^{2}\beta_{k}+\frac{\mu^{\prime}\alpha_{k}}{2}\right)\|x_{k}-x^{*}(y_{k})\|^{2}.$
Now using Assumption 2 and Lemma A.7b), we have that $\mathbb{E}[\|M^{\prime}_{k+1}\|^{2}\mid\mathcal{F}_{k}]$ is bounded by $\mathfrak{c}_{2}\zeta_{2}\left(1+\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2}\right).$ Hence we have
$\mathbb{E}[\|y_{k+1}-y^{*}\|^{2}\mid\mathcal{F}_{k}]\leq\mathfrak{c}_{2}\zeta_{2}\beta_{k}^{2}-\beta_{k}(1-\beta_{k})\|y_{k}-g(x_{k},y_{k})\|^{2}$
$\;\;+\left(1+\frac{2L^{2}\beta_{k}^{2}}{\mu^{\prime}\alpha_{k}}+\mathfrak{c}_{2}\zeta_{2}\beta_{k}^{2}\right)\|y_{k}-y^{*}\|^{2}+\left(L^{2}\beta_{k}+\frac{\mu^{\prime}\alpha_{k}}{2}+\mathfrak{c}_{2}\zeta_{2}\beta_{k}^{2}\right)\|x_{k}-x^{*}(y_{k})\|^{2}.$
Using Assumption 2, we have $\beta_{k}^{2}\leq\alpha_{k}^{3}$. We further have the assumption that $\beta_{k}/\alpha_{k}\leq\mu^{\prime}/(2L^{2})$ (and hence $L^{2}\beta_{k}\leq\mu^{\prime}\alpha_{k}/2$). Taking expectation, we have
$\mathbb{E}\left[\|y_{k+1}-y^{*}\|^{2}\right]$ $\leq\left(1+\Gamma_{2}\alpha_{k}^{2}\right)\mathbb{E}\left[\|y_{k}-y^{*}\|^{2}\right]+(\mu^{\prime}\alpha_{k}+\Gamma_{2}\alpha_{k}^{2})\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$
(14) $\;\;+\Gamma_{2}\alpha_{k}^{2}-\beta_{k}(1-\beta_{k})\mathbb{E}\left[\|y_{k}-g(x_{k},y_{k})\|^{2}\right].$
Here $\Gamma_{2}=2L^{2}/\mu^{\prime}+\mathfrak{c}_{2}\zeta_{2}$. This is our recursive bound on $\mathbb{E}[\|y_{k+1}-y^{*}\|^{2}]$.
A.3 Proof for Lemma 3.3
Proof A.3 (Proof for Lemma 3.3).
In this lemma, we prove that the iterates are bounded. Adding the two intermediate bounds from Lemma 3.2, we get the following.
$\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}+\|y_{k+1}-y^{*}\|^{2}\right]$
$\leq\left(1+(\Gamma_{1}+\Gamma_{2})\alpha_{k}^{2}\right)\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2}\right]+(\Gamma_{1}+\Gamma_{2})\alpha_{k}^{2}.$
Here we have dropped the term $-\beta_{k}(1-\beta_{k})\mathbb{E}\left[\|y_{k}-g(x_{k},y_{k})\|^{2}\right]$ as $10<\beta_{k}<$ 1. Expanding the recursion, we get the following.
$\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2}\right]$
$\leq\prod_{i=0}^{k-1}\left(1+(\Gamma_{1}+\Gamma_{2})\alpha_{i}^{2}\right)\mathbb{E}\left[\|x_{0}-x^{*}(y_{0})\|^{2}+\|y_{0}-y^{*}\|^{2}\right]$
$\;\;+\sum_{i=0}^{k-1}(\Gamma_{1}+\Gamma_{2})\alpha_{i}^{2}\prod_{j=i+1}^{k-1}\left(1+(\Gamma_{1}+\Gamma_{2})\alpha_{j}^{2}\right)$
$\leq e^{\left((\Gamma_{1}+\Gamma_{2})\sum_{i=0}^{\infty}\alpha_{i}^{2}\right)}\left(\mathbb{E}\left[\|x_{0}-x^{*}(y_{0})\|^{2}+\|y_{0}-y^{*}\|^{2}\right]+\sum_{i=0}^{k-1}(\Gamma_{1}+\Gamma_{2})\alpha_{i}^{2}\right).$
Now, we know that the stepsize sequence $\alpha_{k}^{2}$ is summable. Let $D_{1}\coloneqq\sum_{k=0}^{\infty}\alpha_{k}^{2}<\infty$. Then we have
$\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2}\right]$
$\leq e^{D_{1}(\Gamma_{1}+\Gamma_{2})}\left((\Gamma_{1}+\Gamma_{2})D_{1}+\|x_{0}-x^{*}(y_{0})\|^{2}+\|y_{0}-y^{*}\|^{2}\right)$
$\leq\kappa_{1}(\kappa_{1}+\|x_{0}-x^{*}(y_{0})\|^{2}+\|y_{0}-y^{*}\|^{2})\eqqcolon\Gamma_{3}.$
Here $\kappa_{1}=e^{D_{1}(\Gamma_{1}+\Gamma_{2})}$. Hence the iterates $x_{k}$ and $y_{k}$ are bounded in expectation.
A.4 Proof for Lemma 3.4
Proof A.4 (Proof for Lemma 3.4).
This lemma has three parts. First, we show the almost sure convergence of iterates, and then work on mean square bounds for the iterates.
Almost Sure Convergence
We finally prove almost sure convergence for our iterates. We have shown that the iterates satisfy that $\mathbb{E}[\|x_{k}\|^{2}]$ and $\mathbb{E}[\|y_{k}\|^{2}]$ are bounded by some constant (Lemma 3.3) and hence $\|x_{k}\|,\|y_{k}\|<\infty$ almost surely. For a fixed $y$, $f(x,y)$ is contractive in $x$, and hence $x^{*}(y)$ is a globally asymptotically stable equilibrium of the ODE limit (Theorem 12.1 from [2])
$\dot{x}(t)=f(x(t),y)-x(t).$
We have also shown that $x^{*}(y)$ is Lipschitz in $y$ (Lemma 3.1). And hence using Lemma 8.1 from [2], $x_{t}\rightarrow x^{*}(y_{t})$ with probability $1$. Now consider the ODE limit
$\dot{y}(t)=g(x^{*}(y(t)),y(t))-y(t).$
Theorem 12.2 and Remark 3 (mentioned below the theorem) from [2] together imply that $y_{t}$ almost surely converges to the set $\mathcal{Y}^{*}$.
Bound on $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}]$
Applying Lemma 3.3 on the bound from Lemma 3.2, we get
$\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]\leq(1-\mu^{\prime}\alpha_{k})\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]+\Gamma_{1}(1+\Gamma_{3})\alpha_{k}^{2}.$
We get the following bound on expanding the above recursion.
$\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]\leq\|x_{0}-x^{*}(y_{0})\|^{2}\prod_{i=0}^{k-1}(1-\mu^{\prime}\alpha_{i})+\Gamma_{1}(1+\Gamma_{3})\sum_{i=0}^{k-1}\alpha_{i}^{2}\prod_{j=i+1}^{k-1}(1-\mu^{\prime}\alpha_{j}).$
For the first term above, there exists $D_{2}$ such that $1-\mu^{\prime}\alpha_{0}/(k+1)^{\mathfrak{a}}\leq 1-1/(k+D_{2})$. Using Corollary 2.1.2 from [7], we have
$\prod_{i=0}^{k-1}(1-\mu^{\prime}\alpha_{i})\leq\prod_{i=0}^{k-1}\left(1-\frac{1}{i+D_{2}}\right)\leq\frac{D_{2}}{k+D_{2}}.$
For the second term, we use Corollary 2.1.2 from [7] to get the following.
$\sum_{i=0}^{k-1}\alpha_{i}^{2}\prod_{j=i+1}^{k-1}(1-\mu^{\prime}\alpha_{j})\leq\frac{2}{\mu^{\prime}}\alpha_{k}.$
Here we use the fact that $\alpha_{k}$ is of the form $\alpha_{0}/(k+1)^{\mathfrak{a}}$ where $10<\mathfrak{a}<$ 1. Hence we get the following bound.
(15) $\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]\leq\frac{C_{1}}{(k+1)^{\mathfrak{a}}}=D_{3}\alpha_{k}.$
Here $C_{1}=D_{3}\alpha_{0}=\kappa_{2}\left(1+\|x_{0}-x^{*}(y_{0})\|^{2}+\|y_{0}-y^{*}\|^{2}\right)$ where $\kappa_{2}=\max\{\mu^{\prime}D_{2}+2\Gamma_{1}\alpha_{0}\kappa_{1},2\Gamma_{1}\alpha_{0}(1+\kappa_{1}^{2})\}/\mu^{\prime}$. This completes our bound on $\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$.
Bound on summation of $\beta_{i}\mathbb{E}[\|h(y_{i})-y_{i}\|^{2}]$
We first recall that $h(y_{k})=g(x^{*}(y_{k}),y_{k})$. Combining (15) and Lemma 3.3 with the bound in Lemma 3.2,
$\mathbb{E}\left[\|y_{k+1}-y^{*}\|^{2}\right]$
$\leq\mathbb{E}\left[\|y_{k}-y^{*}\|^{2}\right]+(\Gamma_{2}(1+\Gamma_{3})+D_{3}\mu^{\prime})\alpha_{k}^{2}-\beta_{k}(1-\beta_{k})\mathbb{E}\left[\|y_{k}-g(x_{k},y_{k})\|^{2}\right]$
$\leq\mathbb{E}\left[\|y_{k}-y^{*}\|^{2}\right]+(\Gamma_{2}(1+\Gamma_{3})+D_{3}\mu^{\prime}+D_{3}L^{2})\alpha_{k}^{2}-(\beta_{k}/4)\mathbb{E}\left[\|h(y_{k})-y_{k}\|^{2}\right].$
For the second inequality, we use the assumption that $\beta_{k}\leq 1/2$, the fact that $\mathbb{E}[\|h(y_{k})-y_{k}\|^{2}]\leq 2\mathbb{E}[\|g(x_{k},y_{k})-y_{k}\|^{2}]+2L^{2}\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}]$, and that $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}]$ is bounded by $D_{3}\alpha_{k}$.
Hence, we have the following inequality which holds for all $i\geq 0$.
$\frac{\beta_{i}}{4}\mathbb{E}\left[\|h(y_{i})-y_{i}\|^{2}\right]\leq\!\mathbb{E}\left[\|y_{i}-y^{*}\|^{2}\right]-\mathbb{E}\left[\|y_{i+1}-y^{*}\|^{2}\right]+(\Gamma_{2}(1+\Gamma_{3})+D_{3}\mu^{\prime}+D_{3}L^{2})\alpha_{i}^{2}.$
Summing above inequality from $i=0$ to $k$, we get
$\sum_{i=0}^{k}\beta_{i}\mathbb{E}\left[\|h(y_{i})-y_{i}\|^{2}\right]$ $\leq 4\|y_{0}-y^{*}\|^{2}+4(\Gamma_{2}(1+\Gamma_{3})+D_{3}\mu^{\prime}+2D_{3}L^{2})D_{1}\eqqcolon\Gamma_{4},$
where $D_{1}=\sum\alpha_{k}^{2}$ as defined in the proof for Lemma 3.3.
A.5 Proof for Lemma 3.5
Proof A.5 (Proof for Lemma 3.5).
Define $U_{k+1}=(1-\beta_{k})U_{k}+\beta_{k}M^{\prime}_{k+1}$, an averaged error sequence, with $U_{0}=0$. Define $z_{k}=y_{k}-U_{k}$. We first show that $\mathbb{E}[\|z_{k}-y_{k}\|^{2}]$ decays sufficiently fast.
Bound on $\mathbb{E}[\|U_{k}\|^{2}]$
Note that
$U_{k}=\sum_{i=0}^{k-1}\beta_{i}\prod_{j=i+1}^{k-1}(1-\beta_{j})M^{\prime}_{i+1}.$
Using the property that the terms of a martingale difference sequence are orthogonal,
$\mathbb{E}[\|y_{k}-z_{k}\|^{2}]\leq\mathbb{E}\left[\|U_{k}\|^{2}\right]$ $\leq\sum_{i=0}^{k-1}\left(\beta_{i}\prod_{j=i+1}^{k-1}(1-\beta_{j})\right)^{2}\mathbb{E}\left[\|M^{\prime}_{i+1}\|^{2}\right]$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{i=0}^{k-1}\beta_{i}^{2}\prod_{j=i+1}^{k-1}(1-\beta_{j})D_{4}$
(16) $\stackrel{{\scriptstyle(b)}}{{\leq}}2D_{4}\beta_{k}\eqqcolon\Gamma_{5}\beta_{k}.$
Here, inequality (a) follows from Assumption 2 and Lemma 3.3, which together imply that $M^{\prime}_{i+1}$ is bounded in expectation by some constant $D_{4}$. Inequality (b) follows from Corollary 2.1.2 from [7].
Bound on summation of $\beta_{i}\mathbb{E}[\|h(z_{i})-z_{i}\|^{2}]$
We next write $\{z_{k}\}$ as an iteration. We have the iteration $y_{k+1}=(1-\beta_{k})y_{k}+\beta_{k}(g(x_{k},y_{k})+M^{\prime}_{k+1})$. Then using the definition of $z_{k}$, we get
$z_{k+1}+U_{k+1}=(1-\beta_{k})(z_{k}+U_{k})+\beta_{k}g(x_{k},y_{k})+\beta_{k}M^{\prime}_{k+1}.$
Recall that $h(y)=g(x^{*}(y),y)$. This gives us the iteration
$z_{k+1}$ $=(1-\beta_{k})z_{k}+\beta_{k}g(x_{k},y_{k})=z_{k}+\beta_{k}(h(z_{k})-z_{k}+e_{k+1}).$
where $e_{k+1}=g(x_{k},y_{k})-h(y_{k})+h(y_{k})-h(z_{k})$. Next, we write $\|h(z_{i})-z_{i}\|$ in terms of $\|h(y_{i})-y_{i}\|$. For this note the following.
$\|h(z_{i})-z_{i}\|$ $\leq\|h(y_{i})-y_{i}\|+\|h(z_{i})-h(y_{i})\|+\|y_{i}-z_{i}\|\leq\|h(y_{i})-y_{i}\|+2\|y_{i}-z_{i}\|.$
Here, the second inequality follows from the non-expansive nature of function $h(\cdot)$. This gives us the inequality $\|h(z_{i})-z_{i}\|^{2}\leq 2\|h(y_{i})-y_{i}\|^{2}+8\|y_{i}-z_{i}\|^{2}.$ Finally, to obtain a summation similar to Lemma 3.4 c),
$\sum_{i=0}^{k}\beta_{i}\mathbb{E}\left[\|h(z_{i})-z_{i}\|^{2}\right]$ $\leq 2\sum_{i=0}^{k}\beta_{i}\mathbb{E}\left[\|h(y_{i})-y_{i}\|^{2}+4\|y_{i}-z_{i}\|^{2}\right]$
$\leq 2\Gamma_{4}+8\sum_{i=0}^{k}\beta_{i}\mathbb{E}\left[\|y_{i}-z_{i}\|^{2}\right]\leq 2\Gamma_{4}+8\Gamma_{5}\sum_{i=0}^{k}\beta_{i}^{2}.$
Note that $\sum_{i}\beta_{i}^{2}$ is also bounded by $D_{1}$, and hence for $\Gamma_{6}\coloneqq 2\Gamma_{4}+8\Gamma_{5}D_{1}$, we get
(17) $\sum_{i=0}^{k}\beta_{i}\mathbb{E}\left[\|h(z_{i})-z_{i}\|^{2}\right]\leq\Gamma_{6}.$
Bound on $\mathbb{E}[\|h(z_{i})-z_{i}\|^{2}]$ in terms of $\mathbb{E}[\|h(z_{k})-z_{k}\|^{2}]$
For simplicity, we define $p_{i}=h(z_{i})-z_{i}$. We next obtain a bound for $\mathbb{E}[\|p_{i}\|^{2}]$ in terms of $\mathbb{E}[\|p_{k}\|^{2}]$ for all $i\leq k$. Recall that $z_{k+1}=z_{k}+\beta_{k}(h(z_{k})-z_{k}+e_{k+1})=z_{k}+\beta_{k}(p_{k}+e_{k+1})$. We have the following relation.
$\|p_{i+1}\|^{2}$ $=\|p_{i}\|^{2}+\|p_{i+1}-p_{i}\|^{2}+2\langle p_{i+1}-p_{i},p_{i}\rangle$
$=\|p_{i}\|^{2}+\|p_{i+1}-p_{i}\|^{2}+\frac{2}{\beta_{i}}\langle p_{i+1}-p_{i},z_{i+1}-z_{i}-\beta_{i}e_{i+1}\rangle.$
Now,
$2\langle p_{i+1}-p_{i},z_{i+1}-z_{i}\rangle$ $=2\langle h(z_{i+1})-h(z_{i})-z_{i+1}+z_{i},z_{i+1}-z_{i}\rangle$
$=2\langle h(z_{i+1})-h(z_{i}),z_{i+1}-z_{i}\rangle-2\|z_{i+1}-z_{i}\|^{2}$
$=\|h(z_{i+1})-h(z_{i})\|^{2}+\|z_{i+1}-z_{i}\|^{2}-2\|z_{i+1}-z_{i}\|^{2}$
$\;\;-\|h(z_{i+1})-h(z_{i})-z_{i+1}+z_{i}\|^{2}$
$=\|h(z_{i+1})-h(z_{i})\|^{2}-\|z_{i+1}-z_{i}\|^{2}-\|p_{i+1}-p_{i}\|^{2}$
$\leq-\|p_{i+1}-p_{i}\|^{2}.$
Here, the final inequality follows from the non-expansive nature of map $h(\cdot)$. Now,
$\|p_{i+1}\|^{2}$ $\leq\|p_{i}\|^{2}+\|p_{i+1}-p_{i}\|^{2}-\frac{2}{\beta_{i}}\|p_{i+1}-p_{i}\|^{2}-2\langle p_{i+1}-p_{i},e_{i+1}\rangle$
$=\|p_{i}\|^{2}-\frac{1-\beta_{i}}{\beta_{i}}\|p_{i+1}-p_{i}\|^{2}-2\langle p_{i+1}-p_{i},e_{i+1}\rangle.$
Now, note that
$\left\|p_{i+1}-p_{i}+\frac{\beta_{i}}{1-\beta_{i}}e_{i+1}\right\|^{2}=\|p_{i+1}-p_{i}\|^{2}+\left(\frac{\beta_{i}}{1-\beta_{i}}\right)^{2}\|e_{i+1}\|^{2}+\frac{2\beta_{i}}{1-\beta_{i}}\langle p_{i+1}-p_{i},e_{i+1}\rangle.$
Hence
$\|p_{i+1}\|^{2}$ $\leq\|p_{i}\|^{2}-\frac{1-\beta_{i}}{\beta_{i}}\left\|p_{i+1}-p_{i}+\frac{\beta_{i}}{1-\beta_{i}}e_{i+1}\right\|^{2}+\frac{\beta_{i}}{1-\beta_{i}}\|e_{i+1}\|^{2}$
$\leq\|p_{i}\|^{2}+\frac{\beta_{i}}{1-\beta_{i}}\|e_{i+1}\|^{2}\leq\|p_{i}\|^{2}+2\beta_{i}\|e_{i+1}\|^{2}.$
Here, the final inequality follows from the fact that $\beta_{i}\leq 0.5$. Recall that $e_{i+1}=g(x_{i},y_{i})-h(y_{i})+h(y_{i})-h(z_{i})$. Then
$\mathbb{E}\left[\|e_{i+1}\|^{2}\right]$ $\leq 2\mathbb{E}\left[\|g(x_{i},y_{i})-h(y_{i})\|^{2}\right]+2\mathbb{E}\left[\|g(x^{*}(y_{i}),y_{i})-g(x^{*}(z_{i}),z_{i})\|^{2}\right]$
$\leq 2L^{2}\mathbb{E}\left[\|x_{i}-x^{*}(y_{i})\|^{2}\right]+2\mathbb{E}\left[\|y_{i}-z_{i}\|^{2}\right]\leq 2(L^{2}D_{3}+\Gamma_{5})\alpha_{i}.$
Here, the final inequality follows from $\mathbb{E}[\|x_{i}-x^{*}(y_{i})\|^{2}]\leq D_{3}\alpha_{i}$ and $\mathbb{E}[\|y_{i}-z_{i}\|^{2}]\leq\Gamma_{5}\beta_{i}$. We also use the fact that $\beta_{i}\leq\alpha_{i}$ for all $i$. Define $\Gamma_{7}=4(L^{2}D_{3}+\Gamma_{5})$. Hence $\mathbb{E}\left[\|p_{i+1}\|^{2}\right]\leq\mathbb{E}\left[\|p_{i}\|^{2}\right]+\Gamma_{7}\beta_{i}\alpha_{i}.$ This finally gives us the following bound.
$\mathbb{E}\left[\|h(z_{i})-z_{i}\|^{2}\right]\geq\mathbb{E}\left[\|h(z_{k})-z_{k}\|^{2}\right]-\sum_{j=i}^{k-1}\Gamma_{7}\beta_{j}\alpha_{j}.$
A.6 Proof for Theorem 2.1
Proof A.6 (Proof for Theorem 2.1).
We have already shown the almost sure convergence and the bound for $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}]$ in Lemma 3.4. We just need to complete the bound for $\mathbb{E}[\|h(y_{i})-y_{i}\|^{2}]$ to complete the proof. Summing above bound and applying bound from (17), we get
(18) $\mathbb{E}\left[\|h(z_{k})-z_{k}\|^{2}\right]\sum_{i=0}^{k}\beta_{i}\leq\Gamma_{6}+\Gamma_{7}\sum_{i=0}^{k}\beta_{i}\sum_{j=i}^{k-1}\beta_{j}\alpha_{j}.$
To bound the second term, we first note that
[…]
$\stackrel{{\scriptstyle(a)}}{{\leq}}\alpha_{0}\beta_{0}\left(\frac{1}{(i+1)^{\mathfrak{a}+\mathfrak{b}}}+\frac{1}{(\mathfrak{a}+\mathfrak{b}-1)(i+1)^{\mathfrak{a}+\mathfrak{b}-1}}\right)\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{2\alpha_{0}\beta_{0}}{\mathfrak{a}+\mathfrak{b}-1}\frac{1}{(i+1)^{\mathfrak{a}+\mathfrak{b}-1}}.$
Here, inequality (a) and (b) both follow from the fact that $21<\mathfrak{a}+\mathfrak{b}<$ 2, which follows from our assumption that $\alpha_{k}$ and $\beta_{k}$ are both non-summable but square summable sequences. Next,
$\sum_{i=0}^{k}\beta_{i}\sum_{j=i}^{k-1}\beta_{j}\alpha_{j}\leq\frac{2\alpha_{0}\beta_{0}^{2}}{\mathfrak{a}+\mathfrak{b}-1}\sum_{i=0}^{k-1}\frac{1}{(i+1)^{\mathfrak{a}+2\mathfrak{b}-1}}\stackrel{{\scriptstyle(c)}}{{\leq}}\frac{2\alpha_{0}\beta_{0}^{2}}{\mathfrak{a}+\mathfrak{b}-1}\sum_{i=0}^{\infty}\frac{1}{(i+1)^{4\mathfrak{a}-1}}\stackrel{{\scriptstyle(d)}}{{\eqqcolon}}D_{5}.$
Inequality (c) follows from the fact that $2\mathfrak{b}\leq 3\mathfrak{a}$, which follows from our assumption that $\beta_{k}^{2}\leq\alpha_{k}^{3}$. Finally, inequality (d) follows from the fact that $14\mathfrak{a}-1>$ 1, which follows from our assumption that $\mathfrak{a}>0.5$ or $\alpha_{k}$ is square summable.
Hence the right hand side of the inequality (18) is bounded by a constant. Note that $\sum_{i=0}^{k}\beta_{i}\geq(k+1)\beta_{k},$ which follows from our assumption that the sequence $\beta_{k}$ is non-increasing. This implies that
$\mathbb{E}\left[\|h(z_{k})-z_{k}\|^{2}\right]\leq\frac{\Gamma_{6}+\Gamma_{7}D_{5}}{(k+1)\beta_{k}}.$
To complete our proof, we recall that $\|h(y_{k})-y_{k}\|^{2}\leq 2\|h(z_{k})-z_{k}\|^{2}+8\|y_{k}-z_{k}\|^{2}.$ This implies that
(19) $\mathbb{E}\left[\|h(y_{k})-y_{k}\|^{2}\right]$ $\leq\frac{2\Gamma_{6}+2\Gamma_{7}D_{5}}{(k+1)\beta_{k}}+8\Gamma_{5}\beta_{k}\leq\frac{C_{2}}{(k+1)^{1-\mathfrak{b}}},$
where $C_{2}\coloneqq(2\Gamma_{6}+2\Gamma_{7}D_{5})/\beta_{0}+8\Gamma_{5}\beta_{0}$. Here, the first inequality follows from (A.5), and the last inequality follows from the fact that $1/(k+1)^{\mathfrak{b}}\leq 1/(k+1)^{1-\mathfrak{b}}$, which follows from our assumption that $\mathfrak{b}>1/2$. This completes the proof.
A.7 A Technical Lemma
Lemma A.7.
Suppose Assumptions 2 and 2 hold. For all $k\geq 0$ and $y^{*}\in\mathcal{Y}^{*}$, the iterates $\{x_{k},y_{k}\}$ generated by iteration (1) satisfy:
  1. a)
    $\mathbb{E}[\|x^{*}(y_{k+1})-x^{*}(y_{k})\|^{2}\mid\mathcal{F}_{k}]\leq L_{0}^{2}\zeta_{1}\beta_{k}^{2}(1+\|x_{k}\|^{2}+\|y_{k}\|^{2})$ for constant $\zeta_{1}>0$.
  2. b)
    $1+\|x_{k}\|^{2}+\|y_{k}\|^{2}\leq\zeta_{2}(1+\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2})$ for constant $\zeta_{2}>0$.
Proof A.8.
a). Note that
$\|x^{*}(y_{k+1})-x^{*}(y_{k})\|^{2}$ $\leq L_{0}^{2}\|y_{k+1}-y_{k}\|^{2}\leq\beta_{k}^{2}L_{0}^{2}\|g(x_{k},y_{k})-y_{k}+M^{\prime}_{k+1}\|^{2}.$
Using Assumption 2, $\|g(x_{k},y_{k})-y_{k}\|^{2}\leq 2(1+\mathfrak{c}_{1})(1+\|x_{k}\|^{2}+\|y_{k}\|^{2})$. Using Assumption 2, we get $\mathbb{E}[\|M^{\prime}_{k+1}\|^{2}\mid\mathcal{F}_{k}]\leq\mathfrak{c}_{2}(1+\|x_{k}\|^{2}+\|y_{k}\|^{2})$. Combining these bounds completes the proof with $\zeta_{1}=4(1+\mathfrak{c}_{1})+2\mathfrak{c}_{2}$.
b). Note that $\|x_{k}\|^{2}\leq 3\|x_{k}-x^{*}(y_{k})\|^{2}+3\|x^{*}(y_{k})-x^{*}(y^{*})\|^{2}+3\|x^{*}(y^{*})\|^{2}$. Similarly, $\|y_{k}\|^{2}\leq 2\|y_{k}-y^{*}\|^{2}+2\|y^{*}\|^{2}$. Hence
$1+\|x_{k}\|^{2}+\|y_{k}\|^{2}\leq(1+3\|x^{*}(y^{*})\|^{2}+2\|y^{*}\|^{2})+3\|x_{k}-x^{*}(y_{k})\|^{2}+(3L_{0}^{2}+2)\|y_{k}-y^{*}\|^{2}.$
This completes the proof with $\zeta_{2}=\max\{(1+3\|x^{*}(y^{*})\|^{2}+2\|y^{*}\|^{2}),3,(3L_{0}^{2}+2)\}$.
Appendix B Proof for Theorem 2.2
Proof B.1 (Proof for Theorem 2.2).
We define the function $f_{p}(x,y)=\mathcal{P}_{\mathcal{X}}(f(x,y))$ as the projection of function $f(x,y)$ onto the set $\mathcal{X}$. Note that $f_{p}(x,y)$ is also $\mu-$contractive in $x$, i.e., $\|f_{p}(x_{1},y)-f_{p}(x_{2},y)\|\leq\mu\|x_{1}-x_{2}\|$. Similarly, it is $L-$Lipschitz in $y$. This follows from the non-expansive nature of the projection operator. Also, note that $\hat{x}(y)$ has been defined as the fixed point for $f_{p}(\cdot,y)$.
The outline for this proof is the same as the proof for Theorem 2.1.
Lipschitz nature of $\hat{x}(\cdot)$
The map $\hat{x}(\cdot)$ is also Lipschitz with parameter $L_{0}=L/(1-\mu)$. This follows directly from the proof for Lemma 3.1, replacing $f$ with $f_{p}$.
Recursive Bound on $\mathbb{E}[\|x_{k}-\hat{x}(y_{k})\|^{2}]$
Note that
$\|x_{k+1}-\hat{x}(y_{k+1})\|^{2}$ $=\left\|\mathcal{P}_{\mathcal{X}}\big{(}x_{k}+\alpha_{k}(f(x_{k},y_{k})-x_{k}+M_{k+1})\big{)}-\hat{x}(y_{k+1})\right\|^{2}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\left\|\mathcal{P}_{\mathcal{X}}\big{(}x_{k}+\alpha_{k}(f(x_{k},y_{k})-x_{k}+M_{k+1})\big{)}-\mathcal{P}_{\mathcal{X}}(\hat{x}(y_{k+1}))\right\|^{2}$
$\leq\left\|x_{k}+\alpha_{k}(f(x_{k},y_{k})-x_{k}+M_{k+1})-\hat{x}(y_{k+1})\right\|^{2}.$
Here, equality (a) follows from the fact that $\hat{x}(y_{k+1})\in\mathcal{X}$, and the inequality follows from non-expansiveness of the projection operator. Similar to the expansion in (8),
(20a) $\|x_{k+1}-\hat{x}(y_{k+1})\|^{2}=\|x_{k}-\hat{x}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k})\|^{2}$
(20b) $\;\;+\|\alpha_{k}M_{k+1}+\hat{x}(y_{k})-\hat{x}(y_{k+1})\|^{2}$
(20c) $\;\;+2\Big{\langle}x_{k}-\hat{x}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k}),\alpha_{k}M_{k+1}+\hat{x}(y_{k})-\hat{x}(y_{k+1})\Big{\rangle}.$
For the first term (20a), note that
$\|x_{k}-\hat{x}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k})\|^{2}$ $\leq\|x_{k}-\hat{x}(y_{k})\|^{2}+\alpha_{k}^{2}\|f(x_{k},y_{k})-x_{k}\|^{2}$
$\;\;+2\alpha_{k}\langle x_{k}-\hat{x}(y_{k}),f(x_{k},y_{k})-x_{k}\rangle.$
Using Assumption 2, we get $\|f(x_{k},y_{k})-x_{k}\|^{2}\leq 2(1+\mathfrak{c}_{1})(1+\|x_{k}\|^{2}+\|y_{k}\|^{2}).$ For the third term, we first note the property of a projection that $\langle z-\mathcal{P}_{\mathcal{X}}(z),x-\mathcal{P}_{\mathcal{X}}(z)\rangle\leq 0$ for all $z\in\mathbb{R}^{d_{1}},x\in\mathcal{X}$. Hence we get
$\langle f(\hat{x}(y_{k}),y_{k})-\hat{x}(y_{k}),x_{k}-\hat{x}(y_{k})\rangle\leq 0,$
which follows from the fact that $\hat{x}(y_{k})=\mathcal{P}_{\mathcal{X}}(f(\hat{x}(y_{k}),y_{k}))$. This gives us
$\langle x_{k}-\hat{x}(y_{k}),f(x_{k},y_{k})-x_{k}\rangle$ $\leq\Big{\langle}x_{k}-\hat{x}(y_{k}),f(x_{k},y_{k})-x_{k}-(f(\hat{x}(y_{k}),y_{k})-\hat{x}(y_{k}))\Big{\rangle}$
$\leq-\mu^{\prime}\|x_{k}-\hat{x}(y_{k})\|^{2}.$
This follows from the fact that $(x-f(x,y))$ is a $\mu^{\prime}$-strongly monotone function in $x$ for all $y$. This finally gives us
$\|x_{k}-\hat{x}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k})\|^{2}$
(21) $\leq(1-2\mu^{\prime}\alpha_{k})\|x_{k}-\hat{x}(y_{k})\|^{2}+2(1+\mathfrak{c}_{1})\alpha_{k}^{2}(1+\|x_{k}\|^{2}+\|y_{k}\|^{2}).$
The simplification for (20b) is the same as (8b), and hence we do not repeat that here. For (8c),
$2\Big{\langle}x_{k}-\hat{x}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k}),\hat{x}(y_{k})-\hat{x}(y_{k+1})\Big{\rangle}$
$\leq\mu^{\prime}\alpha_{k}\|x_{k}-\hat{x}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k})\|^{2}+\frac{1}{\mu^{\prime}\alpha_{k}}\|\hat{x}(y_{k})-\hat{x}(y_{k+1})\|^{2}.$
The second term is dealt with as before. For the first term, we use (B) as follows
$\mu^{\prime}\alpha_{k}\|x_{k}-\hat{x}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k})\|^{2}$
$\leq\mu^{\prime}\alpha_{k}\|x_{k}-\hat{x}(y_{k})\|^{2}+2(1+\mathfrak{c}_{1})\alpha_{k}^{2}(1+\|x_{k}\|^{2}+\|y_{k}\|^{2}).$
Combining the above results with (A.2) and (A.2), we get the following bound.
$\mathbb{E}\left[\|x_{k+1}-\hat{x}(y_{k+1})\|^{2}\right]$ $\leq\left((1-2\mu^{\prime}\alpha_{k}+\mu^{\prime}\alpha_{k}\right)\mathbb{E}\left[\|x_{k}-\hat{x}(y_{k})\|^{2}\right]$
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!+\left(4(1+\mathfrak{c}_{1})+2\mathfrak{c}_{2}\alpha_{k}^{2}+2L_{0}^{2}\zeta_{1}\beta_{k}^{2}+\frac{L_{0}^{2}\zeta_{1}}{\mu^{\prime}}\frac{\beta_{k}^{2}}{\alpha_{k}}\right)\mathbb{E}\left[1+\|x_{k}\|^{2}+\|y_{k}\|^{2}\right].$
Following the same steps as proof for Lemma 3.2, we get the following bound for $\hat{y}\in\mathcal{\hat{Y}}$. Then we have
$\mathbb{E}\left[\|x_{k+1}-\hat{x}(y_{k+1})\|^{2}\right]$ $\leq(1-\mu^{\prime}\alpha_{k})\mathbb{E}\left[\|x_{k}-\hat{x}(y_{k})\|^{2}\right]$
(22) $\;\;+\Gamma_{1,p}\alpha_{k}^{2}\left(1+\mathbb{E}\left[\|x_{k}-\hat{x}(y_{k})\|^{2}\right]+\mathbb{E}\left[\|y_{k}-\hat{y}\|^{2}\right]\right),$
where $\Gamma_{1,p}\coloneqq\left(4(1+\mathfrak{c}_{1})+2\mathfrak{c}_{2}\zeta_{2}+2L_{0}^{2}\zeta_{1}\zeta_{2}+2L_{0}^{2}\zeta_{1}\zeta_{2}/\mu^{\prime}\right)$. This is our intermediate recursive bound on $\mathbb{E}[\|x_{k+1}-\hat{x}(y_{k+1})\|^{2}]$.
Recursive Bound on $\mathbb{E}[\|y_{k}-\hat{y}\|^{2}]$
This bound follows the exact same proof as Lemma 3.2, with $x^{*}(y)$ replaced with $\hat{x}(y)$, $h(y)$ replaced with $h_{p}(y)=g(\hat{x}(y),y)$ and $y^{*}$ replaced with $\hat{y}$.
$\mathbb{E}\left[\|y_{k+1}-\hat{y}\|^{2}\right]$ $\leq\left(1+\Gamma_{2}\alpha_{k}^{2}\right)\mathbb{E}\left[\|y_{k}-\hat{y}\|^{2}\right]+(\mu^{\prime}\alpha_{k}+\Gamma_{2}\alpha_{k}^{2})\mathbb{E}\left[\|x_{k}-\hat{x}(y_{k})\|^{2}\right]$
(23) $\;\;+\Gamma_{2}\alpha_{k}^{2}-\beta_{k}(1-\beta_{k})\mathbb{E}\left[\|y_{k}-g(x_{k},y_{k})\|^{2}\right].$
Here $\Gamma_{2}=2L^{2}/\mu^{\prime}+\mathfrak{c}_{2}\zeta_{2}$.
Boundedness in Expectation
Adding (B) and (B) as done in Lemma 3.3, we get that the iterates are bounded in expectation. For $\kappa_{1,p}=e^{D_{1}(\Gamma_{1,p}+\Gamma_{2})}$,
$\mathbb{E}\left[\|x_{k}-\hat{x}(y_{k})\|^{2}+\|y_{k}-\hat{y}\|^{2}\right]\leq\kappa_{1,p}(\kappa_{1,p}+\|x_{0}-\hat{x}(y_{0})\|^{2}+\|y_{0}-\hat{y}\|^{2})\eqqcolon\Gamma_{3,p}.$
Almost Sure Convergence
Next, we prove almost sure convergence for our iterates. We have shown that the iterates satisfy that $\mathbb{E}[\|x_{k}\|^{2}]$ and $\mathbb{E}[\|y_{k}\|^{2}]$ are bounded by some constant and hence $\|x_{k}\|,\|y_{k}\|<\infty$ almost surely. We have also shown that $\hat{x}(y)$ is Lipschitz in $y$. For a fixed $y$, the ODE limit of the iteration is $\dot{x}(t)=f(x(t),y)-x(t)+\mathcal{B}_{\mathcal{X}}(x(t)).$ Here $\mathcal{B}_{\mathcal{X}}(x(t))$ is the boundary projection term required to keep $x(t)$ inside $\mathcal{X}$ at all times. Now, we also note that $\hat{x}(y)$ is the unique solution to the variational inequality VI$(\mathcal{X},x-f(x,y))$, given by the solution $x\in\mathcal{X}$ which satisfies $\langle z-x,x-f(x,y)\rangle\geq 0$ for all $z\in\mathcal{X}$. This follows from the fact that $x-f(x,y)$ is strongly monotone in $x$ for all $y$, and Proposition 1.5.8 and Theorem 2.3.3 from [15]. Then Theorem 7 from [13] states that the solutions of the ODE above converge to the solution of the VI, which is $\hat{x}(y)$ in this case. And hence using Lemma 8.1 from [2], $x_{t}\rightarrow\hat{x}(y_{t})$ with probability $1$. Now consider the ODE limit $\dot{y}(t)=g(\hat{x}(y(t)),y(t))-y(t).$ Theorem 12.2 and Remark 3 (mentioned below the theorem) from [2] together imply that $y_{t}$ almost surely converges to the set $\mathcal{\hat{Y}}$.
Finite Time Bounds
The rest of the proof for the finite time bounds is the same as the proof for Theorem 2.1, with the only difference in the constants ($\Gamma_{1}$ and $\Gamma_{3}$ are replaced by $\Gamma_{1,p}$ and $\Gamma_{3,p}$, respectively).
Appendix C Proof for Theorem 4.1
Proof C.1 (Proof for Theorem 4.1).
For this proof, we first give intermediate bounds on $\mathbb{E}[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}]$ and $\mathbb{E}[J(y_{k})]$. For simplicity, define $G(x_{k},y_{k})=g(x_{k},y_{k})-y_{k}$.
Intermediate Bound on $\mathbb{E}[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}]$
We first give a recursive bound on $\mathbb{E}[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}]$. Similar to (8), we have the following.
$\|x_{k+1}-x^{*}(y_{k+1})\|^{2}$
(24a) $=\|x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k})\|^{2}$
(24b) $\;\;+\|\alpha_{k}M_{k+1}+x^{*}(y_{k})-x^{*}(y_{k+1})\|^{2}$
(24c) $\;\;+2\Big{\langle}x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k}),\alpha_{k}M_{k+1}+x^{*}(y_{k})-x^{*}(y_{k+1})\Big{\rangle}.$
The term (24a) is bounded by $(1-\mu^{\prime}\alpha_{k})^{2}\|x_{k}-x^{*}(y_{k})\|^{2}$, where $\mu^{\prime}=1-\mu$ (A.2). We deal with the terms (24b) and (24c) separately.
Term 24b Using Lemma 3.1, we simplify it as follows.
$\|\alpha_{k}M_{k+1}+x^{*}(y_{k})-x^{*}(y_{k+1})\|^{2}$ $\leq 2\alpha_{k}^{2}\|M_{k+1}\|^{2}+2L_{0}^{2}\|y_{k+1}-y_{k}\|^{2}.$
Now, $y_{k+1}-y_{k}=\beta_{k}(G(x_{k},y_{k})-G(x^{*}(y_{k}),y_{k}))-\beta_{k}(\nabla J(y_{k}))+\beta_{k}M^{\prime}_{k+1}$. Hence,
(25) $\|y_{k+1}-y_{k}\|^{2}\leq 3L^{2}\beta_{k}^{2}\|x_{k}-x^{*}(y_{k})\|^{2}+3\beta_{k}^{2}\|\nabla J(y_{k})\|^{2}+3\beta_{k}^{2}\|M^{\prime}_{k+1}\|^{2}.$
Using Assumption 4c, we get
$\mathbb{E}[\|\alpha_{k}M_{k+1}+x^{*}(y_{k})-x^{*}(y_{k+1})\|^{2}\mid\mathcal{F}_{k}]$
$\leq 2\mathfrak{c}_{2}(\alpha_{k}^{2}+3L_{0}^{2}\beta_{k}^{2})+6L_{0}^{2}\beta_{k}^{2}\|\nabla J(y_{k})\|^{2}+6L_{0}^{2}L^{2}\beta_{k}^{2}\|x_{k}-x^{*}(y_{k})\|^{2}.$
Term 24c To simplify (24c), note that
$2\Big{\langle}x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k}),x^{*}(y_{k})-x^{*}(y_{k+1})\Big{\rangle}$
$\leq\frac{\mu^{\prime}\alpha_{k}}{2}\|x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k})\|^{2}+\frac{2}{\mu^{\prime}\alpha_{k}}\|x^{*}(y_{k})-x^{*}(y_{k+1})\|^{2}.$
Using (25) and the fact that $M_{k+1}$ is a martingale difference sequence, we get
$\mathbb{E}\Big{[}x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k}),\alpha_{k}M_{k+1}+x^{*}(y_{k})-x^{*}(y_{k+1})\big{\rangle}\mid\mathcal{F}_{k}\Big{]}$
$\leq\frac{\mu^{\prime}\alpha_{k}}{2}\|x_{k}-x^{*}(y_{k})\|^{2}+\frac{6L_{0}^{2}}{\mu^{\prime}}\frac{\beta_{k}^{2}}{\alpha_{k}}(\mathfrak{c}_{2}+\|\nabla J(y_{k})\|^{2}+L^{2}\|x_{k}-x^{*}(y_{k})\|^{2}).$
Having obtained a bound on the conditional expectation of the three terms in (24), we now combine the bounds after taking expectation. We additionally use our assumption that $\beta_{k}^{2}\leq\alpha_{k}^{3}\leq\alpha_{k}^{2}$.
$\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]\leq D_{6}\alpha_{k}^{2}\mathbb{E}\left[\|\nabla J(y_{k})\|^{2}\right]+D_{7}\alpha_{k}^{2}$
$\;\;+\left(1-2\mu^{\prime}\alpha_{k}+\mu^{\prime 2}\alpha_{k}^{2}+\frac{\mu^{\prime}\alpha_{k}}{2}+6L_{0}^{2}L^{2}\alpha_{k}^{2}+\frac{6L_{0}^{2}L^{2}}{\mu^{\prime}}\alpha_{k}^{2}\right)\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right].$
Here $D_{6}=6L_{0}^{2}(1+1/\mu^{\prime})$ and $D_{7}=2\mathfrak{c}_{2}(1+3L_{0}^{2}+3L_{0}^{2}/\mu^{\prime})$. For appropriate choice of $C_{4}$ in our assumption that $\alpha_{0}\leq C_{4}$, we get $(\mu^{\prime 2}+6L_{0}^{2}L^{2}+6L_{0}^{2}L^{2}/\mu^{\prime})\alpha_{k}^{2}\leq\mu^{\prime}\alpha_{k}/2$.
$\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]$
$\leq(1-\mu^{\prime}\alpha_{k})\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]+D_{6}\alpha_{k}^{2}\mathbb{E}\left[\|\nabla J(y_{k})\|^{2}\right]+D_{7}\alpha_{k}^{2}.$
This gives us
$\mu^{\prime}\alpha_{k}\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]\leq\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]-\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+D_{6}\alpha_{k}^{2}\mathbb{E}\left[\|\nabla J(y_{k})\|^{2}\right]+D_{7}\alpha_{k}^{2}.$
Our intermediate bound for $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}]$ is the following.
$\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]\leq\frac{1}{\mu^{\prime}\alpha_{k}}\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]-\frac{1}{\mu^{\prime}\alpha_{k}}\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]$
(26) $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+(D_{6}/\mu^{\prime})\alpha_{k}\mathbb{E}\left[\|\nabla J(y_{k})\|^{2}\right]+(D_{7}/\mu^{\prime})\alpha_{k}.$
Recursive Bound on $\mathbb{E}[J(y_{k})]$
Note that $\nabla J(\cdot)$ is Lipschitz with parameter $L_{J}\coloneqq L(L_{0}+1)$. To see this, note that
$\|\nabla J(y_{2})-\nabla J(y_{1})\|$ $\leq\|G(x^{*}(y_{2}),y_{2})-G(x^{*}(y_{1}),y_{1})\|$
$\leq L(\|x^{*}(y_{2})-x^{*}(y_{1})\|+\|y_{2}-y_{1}\|)\leq L(L_{0}+1)\|y_{2}-y_{1}\|.$
Then, using the definition of smoothness, we have the following. $J(y_{2})\leq J(y_{1})+\langle\nabla J(y_{1}),y_{2}-y_{1}\rangle+\frac{L_{J}}{2}\|y_{2}-y_{1}\|^{2}.$ Substituting $y_{2}=y_{k+1}$ and $y_{1}=y_{k}$,
(27) $J(y_{k+1})\leq J(y_{k})+\langle\nabla J(y_{k}),y_{k+1}-y_{k}\rangle+\frac{L_{J}}{2}\|y_{k+1}-y_{k}\|^{2}.$
We have already bounded the last term in (25). We simplify the first term as follows.
Term $\langle\nabla J(y_{k}),y_{k+1}-y_{k}\rangle$ Using the simplification above (25), we get
$\langle\nabla J(y_{k}),y_{k+1}-y_{k}\rangle$
$=\beta_{k}\langle\nabla J(y_{k}),G(x_{k},y_{k})-G(x^{*}(y_{k}),y_{k})\rangle-\beta_{k}\|\nabla J(y_{k})\|^{2}+\beta_{k}\langle\nabla J(y_{k}),M^{\prime}_{k+1}\rangle.$
Since $M^{\prime}_{k+1}$ is a martingale difference sequence, we have $\mathbb{E}[\langle\nabla J(y_{k}),M^{\prime}_{k+1}\rangle\mid\mathcal{F}_{k}]=0$. For the first term, we use Cauchy-Schwarz inequality, Young’s inequality, and the Lipschitzness of function $G(\cdot)$ to obtain -
$\beta_{k}\langle\nabla J(y_{k}),G(x_{k},y_{k})-G(x^{*}(y_{k}),y_{k})\rangle\leq\beta_{k}\frac{1}{4}\|\nabla J(y_{k})\|^{2}+\beta_{k}L^{2}\|x_{k}-x^{*}(y_{k})\|^{2}.$
Hence, we get the following.
$\mathbb{E}\left[\langle\nabla J(y_{k}),y_{k+1}-y_{k}\rangle\mid\mathcal{F}_{k}\right]=L^{2}\beta_{k}\|x_{k}-x^{*}(y_{k})\|^{2}-\frac{3}{4}\beta_{k}\|\nabla J(y_{k})\|^{2}.$
We now return to (27). Using our assumption that $\beta_{k}^{2}\leq\beta_{k}$, and taking expectation, we get our intermediate bound on $\mathbb{E}[J(y_{k})]$.
$\mathbb{E}[J(y_{k+1})]$ $\leq\mathbb{E}[J(y_{k})]+\beta_{k}(L^{2}+1.5L_{J}L^{2})\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$
(28) $\;\;+(-0.75\beta_{k}+1.5L_{J}\beta_{k}^{2})\mathbb{E}\left[\|\nabla J(y_{k})\|^{2}\right]+1.5L_{J}\mathfrak{c}_{2}\beta_{k}^{2}.$
Final Bounds
Having obtained intermediate bounds, we substitute the bound for $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}]$ into our bound for $\mathbb{E}[J(y_{k+1})]$. This gives us
$\mathbb{E}[J(y_{k+1})]$ $\leq\mathbb{E}[J(y_{k})]+D_{8}(\beta_{k}/\alpha_{k})\left(\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]-\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]\right)$
$+D_{6}D_{8}\beta_{k}\alpha_{k}\mathbb{E}\left[\|\nabla J(y_{k})\|^{2}\right]+D_{7}D_{8}\beta_{k}\alpha_{k}$
$+(-0.75\beta_{k}+1.5L_{J}\beta_{k}^{2})\mathbb{E}\left[\|\nabla J(y_{k})\|^{2}\right]+1.5L_{J}\mathfrak{c}_{2}\beta_{k}^{2}.$
Here, $D_{8}=(L^{2}+1.5L_{J}L^{2})/\mu^{\prime}$. For appropriate choice of $C_{4}$ in our assumption that $\alpha_{0}\leq C_{4}$, we get $\beta_{k}^{2}\leq\beta_{k}\alpha_{k}$, and $(D_{6}D_{8}+1.5L_{J})\beta_{k}\alpha_{k}\leq 0.5\beta_{k}$.
$\mathbb{E}[J(y_{k+1})]$ $\leq\mathbb{E}[J(y_{k})]-0.25\beta_{k}\mathbb{E}\left[\|\nabla J(y_{k})\|^{2}\right]+(D_{7}D_{8}+1.5L_{J}\mathfrak{c}_{2})\beta_{k}\alpha_{k}$
$\;\;+D_{8}(\beta_{k}/\alpha_{k})\left(\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]-\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]\right).$
This gives us the following relation for all $i\geq 0$.
$\beta_{i}\mathbb{E}\left[\|\nabla J(y_{i})\|^{2}\right]$ $\leq 4\mathbb{E}[J(y_{i})]-4\mathbb{E}[J(y_{i+1})]+4(D_{7}D_{8}+1.5L_{J}\mathfrak{c}_{2})\beta_{i}\alpha_{i}$
$\;\;+4D_{8}(\beta_{i}/\alpha_{i})\left(\mathbb{E}\left[\|x_{i}-x^{*}(y_{i})\|^{2}\right]-\mathbb{E}\left[\|x_{i+1}-x^{*}(y_{i+1})\|^{2}\right]\right).$
Summing over $i=0$ to $k$ gives us
$\sum_{i=0}^{k}\beta_{i}\mathbb{E}\left[\|\nabla J(y_{i})\|^{2}\right]$ $\leq 4J(y_{0})-4\mathbb{E}[J(y_{k+1})]+4(D_{7}D_{8}+1.5L_{J}\mathfrak{c}_{2})\sum_{i=0}^{k}\beta_{i}\alpha_{i}$
(29) $\;\;+4D_{8}\sum_{i=0}^{k}\frac{\beta_{i}}{\alpha_{i}}\left(\mathbb{E}\left[\|x_{i}-x^{*}(y_{i})\|^{2}\right]-\mathbb{E}\left[\|x_{i+1}-x^{*}(y_{i+1})\|^{2}\right]\right).$
We first need to bound the summation on the right-hand side to complete our bound.
$\sum_{i=0}^{k}\frac{\beta_{i}}{\alpha_{i}}\left(\mathbb{E}\left[\|x_{i}-x^{*}(y_{i})\|^{2}\right]-\mathbb{E}\left[\|x_{i+1}-x^{*}(y_{i+1})\|^{2}\right]\right)$
$\leq(\beta_{0}/\alpha_{0})\|x_{0}^{2}-x^{*}(y_{0})\|^{2}-(\beta_{k}/\alpha_{k})\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]$
$\;\;-\sum_{i=1}^{k}\mathbb{E}[\|x_{i}-x^{*}(y_{i})\|^{2}]\left(\frac{\beta_{i-1}}{\alpha_{i-1}}-\frac{\beta_{i}}{\alpha_{i}}\right)\leq(\beta_{0}/\alpha_{0})\|x_{0}^{2}-x^{*}(y_{0})\|^{2}.$
Here, the final inequality follows from the fact that $\beta_{k}/\alpha_{k}$ is a decreasing sequence. Returning to (C), we have the assumption that the sequence $\alpha_{k}\beta_{k}$ is $\mathcal{O}(1/(k+1))$. We also have the assumption that the function $J(\cdot)$ is lower bounded. Hence the right-hand side of (C) is bounded by $D_{9}(1+\log(k+1))$ for some constant $D_{9}>0$. Hence,
$D_{9}(1+\log(k+1))\geq\sum_{i=0}^{k}\beta_{i}\mathbb{E}\left[\|\nabla J(y_{i})\|^{2}\right]$ $\geq\left(\min_{0\leq i\leq k}\mathbb{E}\left[\|\nabla J(y_{i})\|^{2}\right]\right)\sum_{i=0}^{k}\beta_{i}$
$\geq\left(\min_{0\leq i\leq k}\mathbb{E}\left[\|\nabla J(y_{i})\|^{2}\right]\right)k\beta_{k}.$
This completes our proof with
$\min_{0\leq i\leq k}\mathbb{E}\left[\|\nabla J(y_{i})\|^{2}\right]\leq\frac{C_{3}(1+\log(k+1))}{(k+1)^{1-\mathfrak{b}}},$
where $C_{3}=D_{9}/\beta_{0}$. The required $C_{4}$ in the assumption $\alpha_{0}\leq C_{4}$ for the result to hold is $C_{4}=\min\{1,(1-\mu)^{4}/(2+24L^{4}),(1-\mu)^{5}/(18(L^{6}+L^{5}+L^{4}+L^{2}+L))\}$.
Appendix D Proofs from Section 5: Applications
We first prove a lemma that relates strongly monotone and co-coercive operators with contractive and non-expansive mappings, respectively.
Lemma D.1.
Consider function $h:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$. Suppose $\mathfrak{c},\ell>0$. Then,
  1. a).
    If $h(\cdot)$ is $\mathfrak{c}$-strongly monotone and $\ell$-Lipschitz, then function $q(x)=x-(\mathfrak{c}/\ell^{2})h(x)$ is $\sqrt{1-c^{2}/\ell^{2}}$-contractive.
  2. b).
    If $h(\cdot)$ is $\mathfrak{c}$-co-coercive, then function $q(x)=x-2\mathfrak{c}h(x)$ is non-expansive.
Proof D.2.
a). For $x_{1},x_{2}\in\mathbb{R}^{d}$,
$\|q(x_{1})-q(x_{2})\|^{2}$ $=\|x_{1}-x_{2}\|^{2}+\frac{\mathfrak{c}^{2}}{\ell^{4}}\|h(x_{1})-h(x_{2})\|^{2}-2\frac{\mathfrak{c}}{\ell^{2}}\langle h(x_{1})-h(x_{2}),x_{1}-x_{2}\rangle$
$\leq\left(1+\frac{\mathfrak{c}^{2}}{\ell^{4}}\ell^{2}-2\frac{\mathfrak{c}}{\ell^{2}}\mathfrak{c}\right)\|x_{1}-x_{2}\|^{2}=(1-\mathfrak{c}^{2}/\ell^{2})\|x_{1}-x_{2}\|^{2}.$
The inequality follows from the strongly monotone and Lipschitz nature of $h(\cdot)$.
b). For $x_{1},x_{2}\in\mathbb{R}^{d}$,
$\|q(x_{1})-q(x_{2})\|^{2}$ $=\|x_{1}-x_{2}\|^{2}+4\mathfrak{c}^{2}\|h(x_{1})-h(x_{2})\|^{2}-4\mathfrak{c}\langle h(x_{1})-h(x_{2}),x_{1}-x_{2}\rangle$
$\leq\|x_{1}-x_{2}\|^{2}+(4\mathfrak{c}^{2}-4\mathfrak{c}^{2})\|h(x_{1})-h(x_{2})\|^{2}=\|x_{1}-x_{2}\|^{2}.$
The inequality follows from the co-coercive nature of $h(\cdot)$.
D.1 Proof for Corollary 5.1
Proof D.3 (Proof for Corollary 5.1).
Since $H(x,y)$ is $\rho$-strongly-concave in $x$, $-\nabla_{x}H(x,y)$ is $\rho$-strongly monotone [22, Theorem 2.1.9]. Define $f(x,y)=x+(\rho/L^{2})\nabla_{x}H(x,y)$. Then $f(x,y)$ is $\mu-$contractive in $x$ for fixed $y$, where $\mu=\sqrt{1-\rho^{2}/L^{2}}$.
Recall that $\Phi(y)=\max_{x\in\mathbb{R}^{d_{1}}}H(x,y)$. Then Lemma 4.3 from [20] shows that $\nabla\Phi(y)=\nabla_{y}H(x^{*}(y),y)$, and $\nabla\Phi(y)$ is $2L_{0}L$-Lipschitz where $L_{0}=L/(1-\mu)$. Since $H(x,y)$ is convex in $y$, $\Phi(y)=\max_{x}H(x,y)$ is also convex in $y$. This implies that $\nabla\Phi(y)$ is $1/(2LL_{0})$ co-coercive [22, Theorem 2.1.5]. Now, let $g(x,y)=y-\frac{1}{LL_{0}}\nabla_{y}H(x,y)$, and hence $g(x^{*}(y),y)=y-\frac{1}{LL_{0}}\nabla\Phi(y)$. Hence $g(x^{*}(y),y)$ is non-expansive.
The iterates in (3) can then be written as
$x_{k+1}=x_{k}+\alpha_{k}^{\prime}(f(x_{k},y_{k})-x_{k}+M_{k+1})\;\;\text{and}\;\;y_{k+1}=y_{k}+\beta_{k}^{\prime}(g(x_{k},y_{k})-y_{k}+M^{\prime}_{k+1}),$
where $\alpha_{k}^{\prime}=(L^{2}/\rho)\alpha_{k}$ and $\beta_{k}^{\prime}=LL_{0}\beta_{k}$. Then above iteration satisfies Assumptions 2-2. Application of Theorem 2.1 gives us the required result.
D.2 Proof for Corollary 5.2
Proof D.4 (Proof for Corollary 5.2).
The matrix $A_{11}$ is positive definite and hence satisfies $x^{T}A_{11}x\geq\lambda_{A_{11},min}\|x\|^{2}$ and $\|A_{11}x\|\leq\lambda_{A_{11},max}\|x\|$. Here $\lambda_{A_{11},min},\lambda_{A_{11},max}>0$ are the smallest and largest eigenvalues of $A_{11}$, respectively. Hence $f(x,y)=x-\frac{\lambda_{A_{11},min}}{\lambda_{A_{11},max}^{2}}(A_{11}x+A_{12}y-b_{1})$ is $\mu$-contractive in $x$, where $\mu=\sqrt{1-\lambda_{A_{11},min}^{2}/\lambda_{A_{11},max}^{2}}$.
Now, $\Delta=A_{22}-A_{21}A_{11}^{-1}A_{12}$ is non-zero PSD, and hence $\Delta$ is $1/\lambda_{\Delta,max}$-co-coercive [15, Pg. 79]. Then define $g(x,y)=y-(2/\lambda_{\Delta,max})(A_{21}x+A_{22}y-b_{2})$, and hence $g(x^{*}(y),y)$ is non-expansive. Here $x^{*}(y)=A_{11}^{-1}(-A_{12}y+b_{1})$.
Finally, the iterates in (5) can be rewritten as
$x_{k+1}=x_{k}+\alpha_{k}^{\prime}(f(x_{k},y_{k})-x_{k}+M_{k+1})\;\;\text{and}\;\;y_{k+1}=y_{k}+\beta_{k}^{\prime}(g(x_{k},y_{k})-y_{k}+M^{\prime}_{k+1}),$
where $\alpha_{k}^{\prime}=\alpha_{k}(\lambda_{A_{11},max}^{2}/\lambda_{A_{11},min})$ and $\beta_{k}^{\prime}=(\lambda_{\Delta,max}/2)\beta_{k}$. Also, note that
$M_{k+1}$ $=\frac{\lambda_{A_{11},min}}{\lambda_{A_{11},max}^{2}}\left(\tilde{b}_{1}^{(k+1)}-b_{1}-(\tilde{A}_{11}^{(k+1)}-A_{11})x-(\tilde{A}_{12}^{(k+1)}-A_{12})y\right)$
$M^{\prime}_{k+1}$ $=\frac{2}{\lambda_{\Delta,max}}\left(\tilde{b}_{2}^{(k+1)}-b_{2}-(\tilde{A}_{21}^{(k+1)}-A_{21})x-(\tilde{A}_{22}^{(k+1)}-A_{22})y\right)$
satisfy Assumption 2. Theorem 2.1 can then be applied to get the required result.
D.3 Proof for Lemma 5.3
Proof D.5 (Proof for Lemma 5.3).
a). Recall that $x^{*}(\lambda)$ is the unique solution to $\nabla H_{0}(x)=A^{T}\lambda$. Since $H_{0}(\cdot)$ is strongly concave, we have
$\rho\|x^{*}(\lambda_{1})-x^{*}(\lambda_{2})\|^{2}\leq-\langle x^{*}(\lambda_{1})-x^{*}(\lambda_{2}),\nabla H_{0}(x^{*}(\lambda_{1}))-\nabla H_{0}(x^{*}(\lambda_{2}))\rangle$
$=-\langle x^{*}(\lambda_{1})-x^{*}(\lambda_{2}),A^{T}\lambda_{1}-A^{T}\lambda_{2}\rangle=-\langle Ax^{*}(\lambda_{1})-Ax^{*}(\lambda_{2}),\lambda_{1}-\lambda_{2}\rangle,$
for $\lambda_{1},\lambda_{2}\in\mathbb{R}^{d_{2}}$. We have the assumption that $A$ has linearly independent rows, and hence $(AA^{T})^{-1}A$ is the pseudoinverse of matrix $A^{T}$. Hence we have
$(AA^{T})^{-1}A(\nabla H_{0}(x^{*}(\lambda_{1}))-\nabla H_{0}(x^{*}(\lambda_{2})))=\lambda_{1}-\lambda_{2}.$
This implies that
$\|\lambda_{1}-\lambda_{2}\|$ $\leq\|(AA^{T})^{-1}A\|\|\nabla H_{0}(x^{*}(\lambda_{1}))-\nabla H_{0}(x^{*}(\lambda_{2}))\|$
$\leq L\|(AA^{T})^{-1}A\|\|x^{*}(\lambda_{1})-x^{*}(\lambda_{2})\|.$
Hence $\langle-Ax^{*}(\lambda_{1})+Ax^{*}(\lambda_{2}),\lambda_{1}-\lambda_{1}\rangle\geq\rho/(L\|(AA^{T})^{-1}A\|)^{2}\|\lambda_{1}-\lambda_{2}\|^{2}$. This implies that $-Ax^{*}(\lambda)$ is strongly monotone in $\lambda$.
b). Recall that $\hat{x}(\cdot)$ is the unique solution to $\mathcal{P}_{\mathcal{X}}(x+\nabla H_{0}(x)-A^{T}\lambda)=x$. Using [15, Prop. 1.4.2], $\hat{x}(\lambda)$ satisfies: $\left\langle x-\hat{x}(\lambda),\nabla H_{0}(\hat{x}(\lambda))-A^{T}\lambda\right\rangle\leq 0,\forall x\in\mathcal{X},\lambda\in\mathbb{R}^{d_{2}}$. Then we have
$\left\langle\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2}),\nabla H_{0}(\hat{x}(\lambda_{2}))-A^{T}\lambda_{2}\right\rangle$ $\leq 0$
$\left\langle\hat{x}(\lambda_{2})-\hat{x}(\lambda_{1}),\nabla H_{0}(\hat{x}(\lambda_{1}))-A^{T}\lambda_{1}\right\rangle$ $\leq 0$
Combining these equations gives us
$-\left\langle\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2}),\nabla H_{0}(\hat{x}(\lambda_{1}))-\nabla H_{0}(\hat{x}(\lambda_{2}))\right\rangle\leq-\left\langle\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2}),A^{T}\lambda_{1}-A^{T}\lambda_{2}\right\rangle.$
Then, we have the following.
$\rho\|\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2})\|^{2}\leq-\langle\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2}),\nabla H_{0}(\hat{x}(\lambda_{1}))-\nabla H_{0}(\hat{x}(\lambda_{2}))\rangle$
$\leq-\langle\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2}),A^{T}\lambda_{1}-A^{T}\lambda_{2}\rangle=-\langle A\hat{x}(\lambda_{1})-A\hat{x}(\lambda_{2}),\lambda_{1}-\lambda_{2}\rangle,$
Now $\|A\hat{x}(\lambda_{1})-A\hat{x}(\lambda_{2})\|^{2}\leq\|A\|^{2}\|\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2})\|^{2}$, and hence $-A\hat{x}(\lambda)$ is $(\rho/\|A\|^{2})$-co-coercive.
D.4 Proof for Corollary 5.4
Proof D.6 (Proof for Corollary 5.4).
The iterates in (7) can be rewritten as
$x_{k+1}=\mathcal{P}_{\mathcal{X}}\big{(}x_{k}+\alpha_{k}^{\prime}(f(x_{k},\lambda_{k})-x_{k}+M_{k+1})\big{)}\;\;\text{and}\;\;\lambda_{k+1}=\lambda_{k}+\beta_{k}^{\prime}(g(x_{k},\lambda_{k})-\lambda_{k}),$
where $f(x,\lambda)=(\rho/L^{2})(\nabla_{x}H_{0}(x)-A^{T}\lambda)+x,\;g(x,\lambda)=(2\rho/\|A\|^{2})(-Ax+b_{0})+\lambda$ and $\alpha_{k}^{\prime}=(L^{2}/\rho)\alpha_{k},\;\beta_{k}^{\prime}=(\|A\|^{2}/2\rho)\beta_{k}$. Then using Lemma 5.3, the above iteration satisfies assumption 2.1. Assumption 5.3 b) specifies the Slater’s condition, implying existence of $\lambda$ such that $A\hat{x}(\lambda)=b_{0}$ [1, Proposition 27.21]. Application of Theorem 2.2 gives us bounds on $\mathbb{E}[\|g(x_{k},\lambda_{k})-\lambda_{k}\|^{2}]$ which is equal to $(4\rho^{2}/\|A\|^{4})\mathbb{E}[\|Ax-b_{0}\|^{2}]$. This completes our proof for Corollary 5.4.